### v1

# GLProtein: Global-and-Local Structure Aware Protein Representation Learning

Abstract
Proteins are central to biological systems, participating as building blocks across all forms of life. Despite advancements in understanding protein functions through protein sequence analysis, there remains potential for further exploration in integrating protein structural information. We argue that the structural information of proteins is not only limited to their 3D information but also encompasses information from amino acid molecules (local information) to protein-protein structure similarity (global information). To address this, we propose GLProtein, the first framework in protein pre-training that incorporates both global structural similarity and local amino acid details to enhance prediction accuracy and functional insights. GLProtein innovatively combines protein-masked modelling with triplet structure similarity scoring, protein 3D distance encoding and substructure-based amino acid molecule encoding. Experimental results demonstrate that GLProtein outperforms previous methods in several bioinformatics tasks, including predicting protein-protein interaction, contact prediction, and so on.
1 Introduction
Proteins are fundamental to virtually every biological process, serving as the building blocks for cells and organs and acting as catalysts, messengers, and structural elements in all life forms. Understanding the structure and function of proteins is crucial for advances in health guo2017roles ; williams2019plasma ; li2022dietary ; ding2019selective , agriculture rischer2020cellular ; peydayesh2021protein ; davis2024food , and environmental science wheeler2021environmental ; sobhi2023advanced ; zhao2024holistic ; weng2021late , making protein research a cornerstone of biotechnology and medicinal science. Recognizing the critical role of proteins in various scientific fields, many efforts have been made to design computational methods to further understand these crucial molecules sliwoski2014computational ; zhao2020exploring ; li2024tomg . Particularly, protein representation learning, as one significant part, involves capturing the complex features and relationships within proteins in a condensed form that can be utilized for various computational tasks and analyses. It is crucial for enhancing the understanding of protein structures and functions, improving predictive modelling in bioinformatics, facilitating the drug discovery process, and advancing our knowledge of biological systems through interpretable and efficient representations of proteins somnath2021multi ; liu2023generative ; gao2024drugclip .
Figure 1: An illustration on protein representation learning flow. Protein information from local information (inside proteins) to global information (between proteins) can be used as input. This input undergoes encoding by a protein encoder to generate a protein representation across various downstream tasks.
In recent years, the success of language models zhao2024recommender ; hu2023improving ; li2024empowering ; li2024molreflect in natural language processing (NLP) has paved the way for innovative approaches in bioinformatics areas, such as protein modeling xiao2021modeling ; chowdhury2022single , protein generation madani2020progen ; ferruz2022protgpt2 , and protein-protein interaction prediction wang2019high ; ofer2021language . To be specific, by treating protein sequences as linguistic strings, these models have demonstrated remarkable effectiveness in predicting protein function based on sequence data alone. Technically, as shown in Figure 1, protein sequences (e.g., the amino acid sequence ‘MLTAHV…’) are treated as sentences in natural language and amino acids (e.g., ‘M’, ‘L’, and ‘T’) resemble words. Thus, Leveraging the powerful BERT architecture originally developed for natural language, ProtBert elnaggar2021prottrans adeptly adapts the BERT devlin2018bert masked language modelling framework to the field of bioinformatics. This analogy allows ProtBert to employ the technique of predicting randomly masked elements in sequences, thereby learning to identify complex patterns and dependencies among amino acids. Similar to ProtBert, ESM rives2021biological ; verkuil2022language ; hie2022high extends this paradigm by employing a more refined Transformer-based architecture, focusing on capturing the evolutionary relationships and functional dynamics within protein sequences. In other words, most existing protein modelling methods aim to perform protein representation learning by encoding the protein’s sequence information for various downstream applications, such as amino acid contact prediction singh2022spot , protein homology detection kaminski2023plm , protein stability prediction chu2024protein , protein-protein interaction identification wang2019high ; ofer2021language , etc.
Despite the aforementioned successes, most existing protein language modelling methods suffer from intrinsic limitations. Specifically, most of their focuses have primarily been on the amino acid sequence, often neglecting the crucial aspects of protein structure. Proteins possess the ability to fold into diverse 3D shapes, interacting with various proteins and small molecules in biologically significant ways jumper2021highly ; mirdita2022colabfold ; tsaban2022harnessing . Since protein’s structure determines function greslehner2018molecular , utilizing protein 3D structure information effectively is crucial for protein language modelling, in which many studies have demonstrated the potential of pre-training on experimentally determined protein structures hermosilla2022contrastive ; su2023saprot ; wang2022lm ; zhang2022protein . Nevertheless, these methods focus only on the structure within proteins and ignore the global similarities between proteins. We emphasize that the information on protein structure is not only limited to its structure (i.e., conformation) in 3D space but also includes information ranging from local amino acid molecules to the global structural similarity between proteins, as shown in Figure 1. Local information involves the detailed properties and orientations of individual amino acids, which can affect protein stability and biochemical activity renaud2021deeprank . These specifics are vital as they demonstrate how modifications or mutations at the amino acid level can alter the overall structure and functionality of the protein jumper2021highly . Furthermore, protein structure similarities provide information on evolutionary relationships and functional classes, which are crucial for understanding how structurally similar proteins of different species can perform similar or complementary functions within biological systems hamamsy2023protein . For example, as shown in Figure 2, the bacterial ice-binding protein FfIBP and the diatom adhesion protein CaTrailin_4 exhibit no detectable sequence similarity despite their functional similarities zackova2023diatom ; al2023protein . Their predicted structures exhibit a remarkable similarity (TM-Score = 0.6), with both proteins adopting a beta-helical fold comprised of two units linked by an alpha helix. This structural topology is characteristic of ice-binding proteins. Such comparisons are key to predicting the functions of newly discovered proteins based on known structures, thereby enhancing our grasp of complex biological processes and interactions lipman1985rapid ; hamamsy2022tm ; hamamsy2023protein . However, most existing approaches have ineffectively incorporated amino acid molecule information and protein structural similarities into protein representation learning.
(a) FfIBP
(b) CaTrailin_4
(c) Structure Alignment
Figure 2: An example of protein structure similarity. Given the predictive structures of a protein pair: (a) the bacterial ice-binding protein FfIBP and (b) the diatom adhesion protein CaTrailin_4 zackova2023diatom ; al2023protein , (c) is FfIBP (blue) and CaTrailin_4 (green) structure alignment.
To eliminate these limitations, we propose a novel protein pre-training framework GLProtein with Global-and-Local Protein structure information for protein representation learning. Our major contributions are summarized as follows:
  • We introduce a principled approach for capturing protein structural characteristics in a thorough and detailed manner. This approach incorporates a holistic view of protein structure data, encompassing global structural information, protein structure similarities, as well as local structure information such as protein 3D distance encoding and substructure-based molecular encoding. To the best of our knowledge, we are the first to investigate global and local protein structure information in protein language modelling.
  • We propose a novel protein pre-training framework (GLProtein), where protein structure information is incorporated into protein language models for enhancing protein representation learning.
  • The comprehensive experiments demonstrate the effectiveness of the proposed method on a wide range of downstream tasks, which verify the performance superiority of GLProtein.
2 Related Work
In this section, we briefly review related work about protein language modelling as well as protein structure modelling.
Protein Langauge Modeling. As an approach to protein representation learning, protein language modelling is a burgeoning field at the intersection of computational biology and natural language processing (NLP). Inspired by the success of language models in NLP, researchers have adapted these techniques to analyze and predict the properties of protein sequences fan2025computational . Recent advancements have been dominated by the application of transformer-based models, which utilize self-attention mechanisms to capture relationships between amino acids in a sequence. ProtTrans elnaggar2021prottrans and ESM beal2015esm ; verkuil2022language ; hie2022high , trained on large-scale protein databases, have shown remarkable ability in tasks such as protein classification and interaction prediction. Moreover, OntoProtein zhangontoprotein and KeAP zhou2023protein incorporated external biological knowledge to enrich protein representations and enhance performance on various downstream tasks. However, most of these protein language models do not explicitly consider the spatial structure of proteins and structural similarities between proteins like our proposed approach.
Figure 3: Overview of our proposed model, which jointly optimizes global protein similarities and masked protein model with local structure information.
Protein Structure Modelling. The structure of a protein determines its functions. Thus, protein structure modelling has been treated as a reliable way to improve protein representation learning huang2024protein ; alquraishi2021machine ; torrisi2020deep ; cheng2008machine . Some methods use Graph Neural Networks (GNNs) to handle the complex, non-linear relationships inherent in protein structure jha2022prediction ; reau2023deeprank ; xu2024deeprank . Moreover, RGN2 chowdhury2022single utilized a protein language model to learn structural information from unaligned protein sequences. GearNet zhang2022protein focused on geometric pertaining and learned protein features by utilizing spatial relationships between amino acids. SaProt su2023saprot introduced the concept of a "structure-aware vocabulary" to integrate residue tokens with structure tokens. Similar to the knowledge hancing method, PST chen2024endowing enhances protein language models by integrating structural information through graph transformers to incorporate structural data. Unlike these models, we propose global structure learning and local structure learning methods, which could not only integrate protein structure information and amino acid information but also learn the structure similarity between different proteins by using TM-Score hamamsy2023protein .
3 The Proposed Method GLProtein
In this section, we aim to introduce our proposed framework (GLProtein) as a novel solution to learn global and local protein structure information for protein representation learning.
3.1 An Overview of the Proposed Framework
Most existing protein language modelling methods often have limitations, primarily concentrating on linear sequence information while overlooking the essential aspects of protein structure. To address such limitations, we develop GLProtein, a novel protein pre-training framework that incorporates both global and local protein structure information into protein representation learning. More specifically, in order to integrate structure information effectively, a global protein triplet contrastive learning, protein 3D distance encoding, and substructure-based molecular encoding fusion are introduced to empower the protein language model for protein representation learning. The framework of GLProtein, shown in Figure 3, consists of three components: protein language modelling, global structure information modelling, and local structure information modelling. The first component is protein language modelling, which is the backbone of our proposed framework. It employs random masking to the amino acid sequence and treats local structure information as associated structure information, including 3D distance encoding and substructure-based molecular encoding. The second component is global structure information modelling, as input data in this component is still protein sequence. To obtain a protein triplet, we apply positive and negative protein sampling. After that, we use self-supervised contrastive learning to calculate the protein triplet loss and learn the protein similarity with global structure information by taking advantage of the huge amount of unlabeled data. The third component is local structure information modelling, which integrates local protein structure information in the protein decoder, including 3D distance encoding and substructure-based molecular encoding. Next, we will detail each model component.
3.2 Protein Language Modelling
Drawing inspiration from the success of masked language modelling in natural language processing, masked protein modelling conceals part of a protein sequence and tasks a model with predicting the hidden residues devlin2018bert ; elnaggar2021prottrans . As shown in the center part of Figure  3, protein language modelling forms the backbone of our proposed framework, which aims to learn protein representation. We adopt a masking strategy similar to KeAP zhou2023protein , where 20% of the amino acids in a sequence are randomly selected for masking. Each masked amino acid has an 80% probability of being masked for prediction, a 10% chance of being replaced with a random amino acid, and a 10% chance of remaining unchanged. We then integrate protein 3D distance encoding and substructure-based molecular encoding into a protein decoder, in which we will detail in the local structure information modelling component. Suppose that the number of masked amino acids is $M$ and $x_{i}$ denotes the $i$-th amino acid. $x_{\sim i}$ denotes the sequence of amino acids excluding the masked amino acid at position $i$. We leverage a cross-entropy loss $\mathcal{L}_{MLM}$ to estimate masked amino acids. Formally, the masked protein modelling objective can be defined as:
$\mathcal{L}_{MLM}=-\log\sum_{i\in M}P(x_{i}|x_{\sim i};\theta_{E},\theta_{D}),$ (1)
where $\theta_{E}$ and $\theta_{D}$ denote the parameters of the protein sequence encoder and decoder, respectively. We initialize with a pre-trained BERT-like encoder: ProtBert elnaggar2021prottrans . Next, we will introduce the global structure information modelling part.
3.3 Global Structure Information Modelling
In protein science, proteins follow the sequence-structure-function paradigm koehler2023sequence , where the amino acid sequence dictates the protein’s three-dimensional structure, ultimately determining its function. Therefore, effectively utilizing information about protein 3D structures is essential for improving protein language modelling. However, protein structures encompass more than mere 3D spatial configurations; they also include global structural information that reflects similarities among proteins. Understanding these structural similarities is crucial, as they can provide insights into evolutionary relationships and functional roles, allowing for more accurate predictions of protein behaviour and interactions hamamsy2023protein . To address this, we introduce the concept of global structure information, which contains the structure similarities between proteins, by leveraging of the huge amount of self-supervised signals in protein sequences, as shown in the top of the Figure 3. To be specific, given for each input protein sequence, positive and negative protein sampling is designed to get the triplet $(P,P_{pos},P_{neg})$ for capturing protein structure similarity features. Then, the protein triplets are encoded to protein representation for the calculation of the contrastive learning loss. This optimizes the protein sequence encoder by bringing the representation of the input protein $P$ and its positive samples $P_{pos}$ closer together while pushing the representation of $P$ and its negative samples $P_{neg}$ further apart in the representation space.
3.3.1 Positive and Negative Protein Sampling
TM-score (Template Modeling Score) zhang2004scoring ; xu2010significant is a widely used metric in structural biology for assessing the structural similarity between two protein structures. One of its key advantages is the ability to effectively identify similar three-dimensional structures, even among proteins with low sequence similarity. Additionally, the TM-score is more sensitive to global fold similarity than to local structural variations, allowing it to capture essential information about the overall architecture of proteins. Consequently, we utilize the TM-score to measure structural similarity between proteins, focusing on their overall global structure rather than mere sequence identity. Mathematically, the TM-score can be expressed as:
[…]
where $L_{N}$ is the length of the native structure, $L_{r}$ is the length of the aligned residues to the template structures, $d_{i}$ is the distance between the $i$-th pair of residues, and $d_{0}$ is a scaling factor.
We employ a two-pronged approach that utilizes the TM-Vec model hamamsy2023protein to construct a robust set of positive and negative samples for our protein structure similarity analysis. For positive sample selection, we utilize the TM-Vec model to identify the top-K protein sequences that exhibit the highest TM-score values in relation to the template proteins. This approach ensures that our positive samples share significant structural similarity with the template.
In contrast, our negative sampling strategy employs a stochastic selection process followed by structural dissimilarity confirmation. Initially, we randomly select $n$ proteins from our dataset. Subsequently, we employ the TM-Vec model to compute the TM-score between each selected protein and the template protein. Proteins with a TM-score < 0.2 are classified as negative samples, as this threshold indicates a high degree of structural dissimilarity xu2010significant . This method ensures that our negative samples are structurally distinct from the template, providing a clear contrast to the positive samples.
3.3.2 Protein Triplet Modelling
After positive and negative protein sampling, we obtain the triplet $(P,P_{pos},P_{neg})$. Each protein in the triplet is passed to the protein sequence encoder, resulting in the protein representation, i.e., $\mathbf{E}_{p}\in\mathbb{R}^{L_{p}\times D}$, $\mathbf{E}_{p}^{pos}\in\mathbb{R}^{L_{p}\times D}$ and $\mathbf{E}_{p}^{neg}\in\mathbb{R}^{L_{p}\times D}$. $L_{p}$ denotes the length of amino acid sequence and $D$ stands for the feature dimension.
To refine the embedding space by ensuring that proteins with similar structures or functions are clustered closer together than those that are dissimilar, we need to specify an objective function to optimize. Since the task we focus on in this part is contrastive learning, the protein triplet loss is designed. This loss function operates by comparing three entities: anchor protein $P$, positive protein $P_{pos}$ and negative protein $P_{neg}$. Thus, given protein representation triplet $(P,P_{pos},P_{neg})$, the protein triplet loss $\mathcal{L}_{PTL}$ can be defined as:
$\begin{split}\mathcal{L}_{PTL}(P,P_{pos},P_{neg})=max(||\mathbf{E}_{p}-\mathbf{E}^{pos}_{p}||_{2}-||\mathbf{E}_{p}-\mathbf{E}^{neg}_{p}||_{2}+\epsilon,0),\end{split}$ (3)
where $\mathbf{E}_{p}$, $\mathbf{E}^{pos}_{p}$, $\mathbf{E}^{neg}_{p}\in\mathbb{R}^{L_{p}\times D}$ are protein representation of the triplet $(P,P_{pos},P_{neg})$. $\epsilon$ is a margin between positive and negative pairs.
3.4 Local Structure Information Modelling
Complementing global structural information, local protein information offers additional structure-related insights for protein language modelling. While the global structure information modelling component is designed to identify structural similarities across different proteins, the local structure information modelling component zooms in on the specific, intricate features of a protein’s internal structure, providing a more nuanced understanding. More specifically, in this part, we leverage the local structural details of proteins, including protein 3D distance encoding and substructure-based molecular encoding, to enhance the model’s ability to learn and interpret this local configuration effectively, as shown at the bottom of Figure 3. We will first detail the process of protein 3D distance encoding. Then, the integration of embedding of amino acid molecules will be illustrated.
3.4.1 Protein 3D Distance Encoding
In protein representation learning, the integration of protein 3D coordinates holds immense significance for comprehensively capturing proteins’ structural intricacies and functional behaviours. The 3D coordinates provide critical insights into how proteins fold and interact in three-dimensional space, influencing their stability, activity, and specificity liu2022generating ; peng2022pocket2mol ; su2023saprot . By incorporating 3D structural information into representation learning models, the model can capture proteins’ geometric and spatial characteristics, which are often overlooked by sequence-based approaches. We use AlphaFoldDB [^1] https://alphafold.ebi.ac.uk/ as the 3D protein database and integrate the protein 3D distance encoding ying2021transformers to represent protein 3D structural information to ensure rotational and translational invariance. However, proteins are complex macromolecules composed of numerous amino acids, each consisting of multiple atoms. Consequently, the atomic coordinate data for each protein is extensive. To address this challenge, we propose to take advantage of the alpha-carbon ($\alpha$-C) coordinates rather than the entire protein coordinates in protein representation learning. That is because the $\alpha$-C atoms are central to the protein backbone and provide a simplified yet informative representation of the protein’s 3D structure. By capturing the backbone conformation, $\alpha$-C coordinates effectively convey the protein’s overall shape and folding pattern, which are critical for understanding its function. Moreover, leveraging $\alpha$-C coordinates balances capturing essential structural information and maintaining computational efficiency.
Specifically, the coordinates of each $\alpha$-C are processed to represent the current position of the amino acid in 3D space. Then, we encode the Euclidean distance metric to reflect the spatial relation between any pair of amino acids in the 3D space. Mathematically, given each amino acid pair $(i,j)$, we first process their Euclidean distance with the Gaussian Basis Kernel function scholkopf1997comparing ,
$\phi^{k}_{(i,j)}=-\frac{1}{\sqrt{2\pi}|\sigma^{k}|}\exp(-\frac{1}{2}(\frac{\gamma_{(i,j)}||\mathbf{r_{i}}-\mathbf{r_{j}}||+\beta_{(i,j)}-\mu^{k}}{|\sigma^{k}|})^{2}),$ (4)
where $k=1,\dots,K$. $K$ is the number of Gaussian Basis kernels. Then, the 3D distance encoding can be calculated as follows:
$\Phi^{distance}_{(i,j)}=GELU(\bm{\phi}_{(i,j)}\bm{W}^{1}_{D})\bm{W}^{2}_{D},$ (5)
where $\bm{\phi}_{(i,j)}=[\phi^{1}_{(i,j)};\dots;\phi^{K}_{(i,j)}]^{\top}$.  $\bm{W}^{1}_{D}\in\mathbb{R}^{K\times K}$, $\bm{W}^{2}_{D}\in\mathbb{R}^{K\times 1}$ are learnable parameters. $\gamma_{(i,j)},\beta_{(i,j)}$ are learnable scalars indexed by the pair of amino acid types, and $\mu^{k},\sigma^{k}$ are learnable kernel center and learnable scaling factor of the $k$-th Gaussian Basis Kernel. Denote $\Phi^{distance}$ as the matrix form of the 3D distance encoding, whose shape is $n\times n$.
3.4.2 Substructure-based Molecular Encoding
Amino acid molecules, as more detailed information about protein localization, play a crucial role in protein representation learning, as they form the essential building blocks of proteins and provide the foundational data for understanding protein structure and function lieu2020amino ; lopez2024biochemistry . To learn the fine-grained amino acid structure information, we introduce substructure-based molecular encoding to leverage the inherent relationships between molecule motifs and substructural features in amino acid molecules. In practice, we utilize the mol2vec jaeger2018mol2vec to process and derive representations for all amino acid molecules to obtain fine-grained molecular structure information. Then we construct substructure-based molecular encoding as fine-grained protein sequence representation, which is the other part of local structure information modelling component. For protein $P$, we have
$\mathbf{E}_{a}(P)=\text{Concat}(\mathbf{e}_{x_{1}},\mathbf{e}_{x_{2}},\dots,\mathbf{e}_{x_{i}},\dots,\mathbf{e}_{x_{L}}),\text{where}~{}\mathbf{e}_{x_{i}}\in\mathbb{R}^{1\times d},$ (6)
$L$ is the length of the protein sequence, $\mathbf{e}_{x_{i}}$ is the $i$-th amino acid molecule embedding, and $d$ stands for the feature dimension of the amino acid molecule.
3.5 Model Training
In this part, we will first detail the protein decoder process, which combines protein language modelling and local structure information modelling components. Finally, the pre-training objective of the whole framework will be stated.
3.5.1 Protein Decoder
As shown in Figure 3, to enhance the learning of protein representation, we introduce a protein decoder that enables each protein sequence representation to iteratively query and extract relevant information from its associated local structure, including protein 3D distance encoding and substructure-based molecular encoding. This process facilitates the integration of critical structural details, aiding in the accurate reconstruction of masked amino acids. Specifically, the decoder treats protein representation $\mathbf{E}_{p}$ as a query, while the substructure-based molecular encodings $\mathbf{E}_{a}$ are attended to as keys and values and protein 3D distance encoding $\Phi^{distance}$ is attended to as attention bias. Taking the $i$-th layer as an example, the inputs to the protein decoder include $\mathbf{E}_{p}^{i}$, $\Phi^{distance}$ and $\mathbf{E}_{a}$. The substructure-based molecular encoding $\mathbf{E}_{a}$ is firstly queries by $\mathbf{E}_{p}^{i}$ as the key and value:
$Q_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})W_{Q}^{i},~{}K_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{K}^{i},~{}V_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{V}^{i},$ (7)
where $W_{Q}^{i}$,$W_{K}^{i}$,$W_{V}^{i}$ are learnable matrices. Norm stands for the layer normalization ba2016layer .
Then, Attention vaswani2017attention is applied to $\{Q_{p}^{i},K_{a}^{i},V_{a}^{i}\}$, where the representation of protein sequence extract helpful, relevant information from the substructure-based molecular encoding. The obtained representation $o_{p}^{i}$ stores the helpful structure information for restoring missing amino aids. We then add up $o_{p}^{i}$ and $\mathbf{E}_{p}^{i}$ to integrate information, resulting in the representation $\hat{\mathbf{E}}_{p}^{i}$ as follows:
$o_{p}^{i}=\text{Attention}(Q_{p}^{i},K_{a}^{i},V_{a}^{i},\Phi^{distance}),~{}~{}\hat{\mathbf{E}}_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})+o_{p}^{i}.$ (8)
The resulting representation $\hat{\mathbf{E}}_{p}^{i}$ integrates the helpful, relevant structure information that benefits the restoration of missing amino acids. We finally forward $\hat{\mathbf{E}}_{p}^{i}$ through a residual multi-layer perceptron to obtain the output representation of the $i$-th block, which also serves as the input to the $(i+1)$-th block.
Table 1: Comparisons on amino acid contact prediction. seq signifies the distance, measured in terms of amino acid units, between two selected amino acids. P@L, P@L/2, P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively. The best results are bolded, and the second-best results are underlined.
$126\leq seq<$ 12 $2412\leq seq<$ 24 $24\leq seq$
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
LSTM 0.26(±0.02) 0.36(±0.01) 0.49(±0.03) 0.20(±0.02) 0.26(±0.02) 0.34(±0.03) 0.20(±0.01) 0.23(±0.02) 0.27(±0.02)
ResNet 0.25(±0.02) 0.34(±0.02) 0.46(±0.02) 0.28(±0.01) 0.25(±0.01) 0.35(±0.03) 0.10(±0.03) 0.13(±0.02) 0.17(±0.03)
Transformer 0.28(±0.03) 0.35(±0.01) 0.46(±0.02) 0.19(±0.02) 0.25(±0.02) 0.33(±0.01) 0.17(±0.02) 0.20(±0.02) 0.24(±0.02)
ProtBert 0.30(±0.03) 0.40(±0.02) 0.52(±0.02) 0.27(±0.03) 0.35(±0.02) 0.47(±0.01) 0.20(±0.01) 0.26(±0.02) 0.34(±0.01)
OntoProtein 0.37(±0.02) 0.46(±0.01) 0.57(±0.03) 0.32(±0.01) 0.40(±0.02) 0.50(±0.02) 0.24(±0.03) 0.31(±0.01) 0.39(±0.03)
LM-GVP 0.35(±0.02) 0.42(±0.02) 0.49(±0.02) 0.33(±0.03) 0.43(±0.02) 0.51(±0.03) 0.26(±0.02) 0.37(±0.02) 0.43(±0.03)
GearNet 0.39(±0.02) 0.46(±0.02) 0.57(±0.02) 0.36(±0.03) 0.44(±0.02) 0.55(±0.03) 0.29(±0.02) 0.37(±0.01) 0.45(±0.02)
SaProt 0.41(±0.02) 0.39(±0.03) 0.42(±0.02) 0.38(±0.01) 0.37(±0.01) 0.41(±0.01) 0.24(±0.02) 0.27(±0.03) 0.37(±0.02)
KeAP 0.41(±0.04) 0.52(±0.02) 0.62(±0.03) 0.36(±0.01) 0.45(±0.01) 0.57(±0.01) 0.29(±0.02) 0.37(±0.03) 0.46(±0.02)
ESM-2 0.42(±0.02) 0.49(±0.03) 0.63(±0.01) 0.37(±0.01) 0.43(±0.01) 0.57(±0.02) 0.30(±0.02) 0.38(±0.03) 0.46(±0.02)
GLProtein 0.45(±0.02) 0.55(±0.02) 0.66(±0.01) 0.39(±0.03) 0.48(±0.01) 0.58(±0.02) 0.31(±0.02) 0.40(±0.01) 0.47(±0.02)
Figure 4: An example of amino acid contacts (top-L predictions for ProteinNet alquraishi2019proteinnet test example TBM-hard#T0912). Raw contact probabilities are shown below the diagonal, top L contacts are shown above the diagonal (blue: true positives, red: false positives, grey: ground-truth contacts).
By training with both protein 3D distance encoding and substructure-based molecular encoding, the model can not only go beyond simply memorizing sequence patterns but also actively decode the underlying structural and chemical principles that govern protein functionality.
3.5.2 Pre-training Objective
To estimate the model parameters of GLProtein, we adopt the masked protein modelling object and global protein triplet objective to construct the overall model. We jointly optimize the overall objective as follows:
$\mathcal{L}=\mathcal{L}_{MLM}+\alpha\mathcal{L}_{PTL},$ (9)
where $\alpha$ is the hyper-parameter.
Figure 5: Results on TAPE Benchmark encompass various evaluations. SS is a secondary structure task that is evaluated in CB315. We report medium- and long-range results using P@L/2 metrics in contact prediction task. In fluorescence and stability prediction tasks, we use Spearman’s $\rho$ metric for evaluation. We also provide a related table in Appendix 7.1.
4 Experiments
In this section, we evaluate the generalization ability of the learned protein representation by fine-tuning the pre-trained model across a diverse array of downstream applications, including amino acid contact prediction, protein homology detection, protein stability prediction, protein-protein interaction identification, protein-protein binding affinity prediction and semantic similarity inference.
4.1 Dataset and Implementation Details
ProteinKG25  zhangontoprotein offers a comprehensive knowledge graph comprising approximately five million triplets. This graph includes nearly 600k proteins, 50k attribute terms and 31 relation terms. However, note that we only use the protein sequence data in this dataset. Our model does not use the attribute and relation terms extracted from Gene Ontology. We used modified pre-training data from KeAP zhou2023protein , which removed overlapping data between ProteinKG25 and downstream datasets. Additionally, we use AlphaFoldDB to obtain the protein 3D coordinate datasets.
We conducted some experiments and compared GLProtein with baselines regarding pre-training and inference time in contact prediction tasks, as shown in Appendix Table 4. GLProtein outperforms baselines in multiple downstream tasks with similar parameters. During pre-training, GLProtein is trained for 300k steps using a learning rate of 1e-5, weight decay of 0.01 over four GPUs (NVIDIA A6000, 48G Memory each). For the amino acid contact prediction and protein-protein interaction task, we randomly selected five random seeds to fine-tune our model and the baseline model separately and report the results. For full implementation details, refer to the provided code.
4.2 Downstream Tasks
4.2.1 Amino Acid Contact and Secondary Structure Prediction
Overview of amino acid contact prediction. Amino acid contact prediction is a critical task in computational biology, aiming to identify pairs of amino acids within a protein that are in close spatial proximity. Given an input protein sequence, our model predicts whether pairs of amino acids from the same sequence are in contact. The model accomplishes this by generating a probability contact matrix for each input protein. We tested the model on the dataset collected and organized by ProteinNet alquraishi2019proteinnet and TAPE rao2019evaluating .
Overview of secondary structure prediction. Secondary structure is a fundamental aspect of computational biology, aiming to determine the local structures of protein segments. this task is a sequence-to-sequence task where each input protein is mapped to a type of local structure. We report accuracy on a per-amino acid basis on the CB513 dataset cuff1999evaluation .
Baselines. We evaluate our model compared with ten baselines. Specifically, we employed variations of LSTM hochreiter1997long , ResNet he2016deep and Transformer vaswani2017attention proposed by the TAPE benchmark rao2019evaluating . ProtBert elnaggar2021prottrans is a BERT-like model pre-trained on UniRef100 suzek2007uniref ; suzek2015uniref . ESM-2 rives2021biological ; verkuil2022language ; hie2022high feature a transformer architecture pre-trained on the representative sequences from UniRef50 suzek2007uniref ; suzek2015uniref . OntoProtein zhangontoprotein and KeAP zhou2023protein are the most recent knowledge-based pre-training methodologies. SaProt su2023saprot is the most recent structure-based protein language model. LM-GVP wang2022lm and GearNet zhang2022protein are famous geometric methods for protein representation learning.
Results. Table  1 shows the experimental results of amino acid contact prediction. Specifically, we notice that our model GLProtein consistently outperforms other models in short- ($126\leq seq<$ 12), medium- ($2412\leq seq<$ 24) and long-range ($seq>24$) contact predictions. Notably, our model demonstrates better performance compared to SaProt, which is also a structure-based language model. We also randomly selected a protein from the contact test dataset for visual analysis. As shown in Figure 4, the left is our GLProtein’s result of amino acid contacts. The right three are the contact maps of three baseline models, including KeAP, ESM-2 and ProtBERT. Figure 4 shows more visually that GLProtein’s prediction on the task of contact prediction is closer to labels, i.e., better performance on long-range contact prediction. For the secondary structure (SS-Q3 and SS-Q8), as shown in Figure 5, GLProtein outperforms other baselines in SS-Q8 task and shows competitive performance with ProtBERT, OntoProtein and KeAP in SS-Q3 task. Considering the approaches taken by Saprot, LM-GVP, and GearNet, which also incorporate protein structural information, the evident performance superiority of GLProtein over these methods indicates that it offers a more effective option for structure-based protein representation learning. We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure.
4.2.2 Protein-Protein Interaction
Overview. Protein-protein interaction (PPI) is fundamental to virtually all biological processes and pathways in living organisms. It refers to the physical contact between two or more amino acid sequences. In this paper, we only focus on two-protein cases where a pair of protein sequences serve as the inputs. The objective is to accurately predict the specific types of interactions that occur between each pair of proteins.
In our experiments, we focus on predicting 7 interaction types between protein pairs, namely reaction, binding, post-translational modifications, activation, inhibition, catalysis, and expression. The challenge of PPI prediction is approached as a multi-label classification problem. We conducted our experiments using three datasets: SHS27K chen2019multifaceted , SHS148K chen2019multifaceted and STRING lv2021learning . Both SHS27K and SHS148K are considered subsets of STRING, with proteins excluded if they have fewer than 50 amino acids or exhibit 40% or higher sequence identity. We followed OntoProtein’s setting to generate test sets and employed Breadth-First Search (BFS) and Depth-First Search (DFS) techniques across these datasets. The F1 score is utilized as the primary metric for evaluating performance.
Baselines. Following OntoProtein zhangontoprotein and KeAPzhou2023protein , we have expanded our baseline models to include four additional methods: DPPI hashemifar2018predicting , DNN-PPI li2018deep , PIPR chen2019multifaceted and GNN-PPI lv2021learning . These are incorporated alongside existing baselines such as ProtBert, ESM-2, OntoProtein, KeAP, SaProt, LM-GVP and GearNet, providing a comprehensive set of comparisons in our analysis.
Table 2: Protein-Protein Interaction Prediction Results. Breath-First Search (BFS) and Depth-First Search (DFS) are strategies that split the training and testing PPI datasets. The best results are bolded, and the second-best results are underlined.
SHS27k SHS148k STRING
Methods BFS DFS BFS DFS BFS DFS
DPPI 40.27(±0.74) 44.86(±0.87) 51.26(±0.66) 51.43(±0.94) 55.79(±0.81) 64.72(±0.94)
DNN-PPI 47.97(±0.94) 52.85(±0.91) 55.90(±0.67) 57.82(±0.78) 52.74(±0.89) 62.99(±0.93)
PIPR 43.67(±0.99) 56.76(±0.82) 60.10(±0.85) 61.83(±0.94) 53.65(±0.88) 66.46(±0.92)
GNN-PPI 62.47(±0.65) 73.19(±0.89) 71.01(±0.92) 81.54(±0.87) 75.34(±0.82) 90.01(±0.78)
ProtBert 68.44(±0.78) 72.36(±0.85) 70.06(±0.88) 77.46(±0.62) 66.08(±0.91) 86.45(±0.82)
OntoProtein 71.37(±0.84) 76.28(±0.77) 74.60(±0.56) 76.33(±0.69) 75.64(±0.91) 90.23(±0.79)
KeAP 78.51(±0.95) 78.84(±0.85) 74.26(±0.89) 81.99(±0.92) 80.08(±0.79) 88.47(±0.71)
[…]
GearNet 85.46(±0.61) 82.73(±0.69) 80.02(±1.26) 82.28(±0.93) 85.55(±0.74) 88.03(±0.51)
ESM-2 94.01(±0.77) 87.32(±0.97) 91.46(±0.63) 85.24(±0.46) 88.13(±0.71) 85.53(±0.55)
SaProt 91.18(±0.73) 88.85(±1.04) 90.75(±0.91) 80.67(±0.90) 88.23(±0.81) 88.90(±0.74)
GLProtein 96.32(±0.86) 91.23(±0.92) 93.78(±0.77) 86.14(±0.69) 89.41(±0.66) 91.35(±0.89)
Results. As shown in Table 2, the results clearly indicate that our method consistently outperforms all other methods, including the structure-based protein language model SaProt, across all datasets and both BFS and DFS evaluation metrics. The observed decline in performance can be linked to the growing amount of fine-tuning data, transitioning from SHS27k to STRING, which diminished the influence of pre-training. We believe that the structural similarities between proteins identified during the pre-training step enable GLProtein to excel in the PPI task, resulting in its outstanding performance.
4.2.3 Homology Detection, Fluorescence and Stability Prediction
Overview of homology detection. The task of predicting remote homology in proteins can be viewed as a classification problem at the molecular level. The objective is to input a protein sequence into the homology detection model, which then identifies the correct types of protein fold. In our paper, this presents a significant challenge with 1,195 distinct protein folds to classify. We utilize data sources from  hou2018deepsf and present the average accuracy achieved on the fold-level heldout set.
Overview of fluorescence prediction. In the realm of protein science, fluorescence prediction is a vital task that involves estimating the fluorescence properties of proteins. This is a regression task where each input protein is mapped to a label measuring the most extreme circumstances in which the protein maintains its fold above a concentration threshold. We use the data from  rocklin2017global and use Spearman’s rank correlation coefficient as the metric.
Overview of stability prediction. Stability prediction involves estimating the resilience of a protein’s structure under various environmental conditions, a critical factor in understanding its functional efficacy and therapeutic potential. This regression task focuses on predicting the intrinsic stability of proteins, which is essential for assessing their capacity to preserve their structural integrity under severe conditions. To assess the effectiveness of our model, we measure its performance using Spearman’s rank correlation coefficient across the entire test set rocklin2017global .
Baselines. As shown in Figure 5, we included ten protein model as baselines.
Results. As for fluorescence prediction, Figure 5 shows our model has the most competitive performance compared to Transformer and KeAP. In the domain of stability prediction, our model again shows the highest performance with a score of 0.81. This is significantly higher compared to other models, indicating its potential utility in applications like drug design and protein engineering, where stability is paramount.
4.2.4 Protein Function Prediction
Figure 6: Comparisons on semantic similarity inference. The best results are bolded, and the second-best results are underlined.
Methods MF BP CC Avg
ESM-2 0.31 0.42 0.28 0.34
ProtBert 0.41 0.35 0.36 0.37
OntoProtien 0.41 0.36 0.36 0.38
SaProt 0.40 0.40 0.39 0.40
KeAP 0.40 0.40 0.40 0.40
GLProtein 0.41 0.40 0.39 0.40
Overview. Protein function prediction aims to assign biological or biochemical roles to proteins, and we also regard this task as a sequence classification task. Following KeAP zhou2023protein , we divide protein attributes into three groups: molecular function (MF), biological process (BP) and cellular component (CC), and report the Spearman’s rank correlation scores for each group.
Baselines. We evaluate our model compared with five baselines, including ESM-2, ProtBERT, OntoProtein, SaProt and KeAP.
Results. Table 6 assesses the performance of various computational models in prediction protein functions in three categories: MF, BP and CC. Additionally, an average score (Avg) is calculated for each method to provide a holistic view of performance across all categories. These models all show a balanced performance in three groups. It is worth noting that our model does not use any protein attribute-related knowledge and is comparable to OntoProtein and KeAP, which do. It also demonstrates the superiority of our approach.
We also evaluate our GLProtein with 6 baseline models on protein-protein binding affinity estimation task, which is detailed in Appendix 7.2.
5 Failure Discussion
We have observed that GLProtein underperforms on certain individual tasks. For instance, in the protein-protein binding affinity prediction task, ESM-2 surpasses GLProtein. This task focuses on predicting changes in binding affinity resulting from protein mutations. We believe that GLProtein’s limited performance is attributed to its lack of mutation information, whereas ESM-2 incorporates multiple sequence alignment (MSA) data during training, which includes mutation insights. Similarly, in the Fluorescence task, GLProtein does not demonstrate significant improvement when tasked with distinguishing closely related proteins. We hypothesize that while GLProtein effectively learns structural similarities among different proteins during pre-training, it excels at identifying differences between dissimilar structures but struggles to differentiate between similar ones. We plan to further investigate these issues in our future research.
6 Conclusion and Future Work
In this work, we propose GLProtein, a general protein pre-training framework with global and local protein structure information. By integrating structure similarities between proteins, protein 3D distance encoding and amino acid molecule embeddings, GLProtien outperforms the previous protein representation learning model on most downstream applications, demonstrating the performance superiority of GLProtein. In the future, we aim to further enhance GLProtein’s capabilities by exploring novel avenues for incorporating multi-modal data sources, refining the model’s interpretability, and extending its applicability to a wider array of biological contexts.

### v2

@@
 GLProtein innovatively combines protein-masked modelling with triplet structure similarity scoring, protein 3D distance encoding and substructure-based amino acid molecule encoding.
-Experimental results demonstrate that GLProtein outperforms previous methods in several bioinformatics tasks, including predicting protein-protein interaction, contact prediction, and so on.
+Experimental results demonstrate that GLProtein outperforms previous methods in several bioinformatics tasks, including predicting protein-protein interactions, contact prediction, and so on.
+GLProtein: Global-and-Local Structure Aware Protein Representation Learning
+Yunqing Liu     Wenqi Fanthanks: Corresponding author† ‡     Xiao-Yong Wei† §     Qing Li Department of Computing (COMP), The Hong Kong Polytechnic University Department of Management and Marketing (MM), The Hong Kong Polytechnic University §Department of Computer Science, Sichuan University yunqing617.liu@connect.polyu.hk wenqifan03@gmail.com, {cs007.wei, csqli}@comp.polyu.edu.hk
 1 Introduction
 Proteins are fundamental to virtually every biological process, serving as the building blocks for cells and organs and acting as catalysts, messengers, and structural elements in all life forms.
-Understanding the structure and function of proteins is crucial for advances in health guo2017roles ; williams2019plasma ; li2022dietary ; ding2019selective , agriculture rischer2020cellular ; peydayesh2021protein ; davis2024food , and environmental science wheeler2021environmental ; sobhi2023advanced ; zhao2024holistic ; weng2021late , making protein research a cornerstone of biotechnology and medicinal science.
-Recognizing the critical role of proteins in various scientific fields, many efforts have been made to design computational methods to further understand these crucial molecules sliwoski2014computational ; zhao2020exploring ; li2024tomg .
+Understanding the structure and function of proteins is crucial for advances in health, agriculture, and environmental science, making protein research a cornerstone of biotechnology and medicinal science Ding et al.
+(2019); Davis et al.
+(2024); Zhao et al.
+(2024).
+Recognizing the critical role of proteins in various scientific fields, many efforts have been made to design computational methods to further understand these crucial molecules Weng et al.
+(2021); Zhao et al.
+(2020).
 Particularly, protein representation learning, as one significant part, involves capturing the complex features and relationships within proteins in a condensed form that can be utilized for various computational tasks and analyses.
-It is crucial for enhancing the understanding of protein structures and functions, improving predictive modelling in bioinformatics, facilitating the drug discovery process, and advancing our knowledge of biological systems through interpretable and efficient representations of proteins somnath2021multi ; liu2023generative ; gao2024drugclip .
+It is crucial for enhancing the understanding of protein structures and functions, improving predictive modelling in bioinformatics, facilitating the drug discovery process, and advancing our knowledge of biological systems through interpretable and efficient representations of proteins Somnath et al.
+(2021); Liu et al.
+(2023); Gao et al.
+(2024).
 Figure 1: An illustration on protein representation learning flow.
@@
 This input undergoes encoding by a protein encoder to generate a protein representation across various downstream tasks.
-In recent years, the success of language models zhao2024recommender ; hu2023improving ; li2024empowering ; li2024molreflect in natural language processing (NLP) has paved the way for innovative approaches in bioinformatics areas, such as protein modeling xiao2021modeling ; chowdhury2022single , protein generation madani2020progen ; ferruz2022protgpt2 , and protein-protein interaction prediction wang2019high ; ofer2021language .
+In recent years, the success of language models in natural language processing (NLP) has paved the way for innovative approaches in bioinformatics areas, such as protein modeling Xiao et al.
+(2021); Chowdhury et al.
+(2022), protein generation Madani et al.
+(2020); Ferruz et al.
+(2022), and protein-protein interaction prediction Wang et al.
+(2019); Ofer et al.
+(2021).
 To be specific, by treating protein sequences as linguistic strings, these models have demonstrated remarkable effectiveness in predicting protein function based on sequence data alone.
 Technically, as shown in Figure 1, protein sequences (e.g., the amino acid sequence ‘MLTAHV…’) are treated as sentences in natural language and amino acids (e.g., ‘M’, ‘L’, and ‘T’) resemble words.
-Thus, Leveraging the powerful BERT architecture originally developed for natural language, ProtBert elnaggar2021prottrans adeptly adapts the BERT devlin2018bert masked language modelling framework to the field of bioinformatics.
+Thus, Leveraging the powerful BERT architecture originally developed for natural language, ProtBert Elnaggar et al.
+(2021) adeptly adapts the BERT Devlin et al.
+(2018) masked language modelling framework to the field of bioinformatics.
 This analogy allows ProtBert to employ the technique of predicting randomly masked elements in sequences, thereby learning to identify complex patterns and dependencies among amino acids.
-Similar to ProtBert, ESM rives2021biological ; verkuil2022language ; hie2022high extends this paradigm by employing a more refined Transformer-based architecture, focusing on capturing the evolutionary relationships and functional dynamics within protein sequences.
-In other words, most existing protein modelling methods aim to perform protein representation learning by encoding the protein’s sequence information for various downstream applications, such as amino acid contact prediction singh2022spot , protein homology detection kaminski2023plm , protein stability prediction chu2024protein , protein-protein interaction identification wang2019high ; ofer2021language , etc.
+Similar to ProtBert, ESM Rives et al.
+(2021); Verkuil et al.
+(2022); Hie et al.
+(2022) extends this paradigm by employing a more refined Transformer-based architecture, focusing on capturing the evolutionary relationships and functional dynamics within protein sequences.
+In other words, most existing protein modelling methods aim to perform protein representation learning by encoding the protein’s sequence information for various downstream applications, such as amino acid contact prediction Singh et al.
+(2022), protein homology detection Kaminski et al.
+(2023), protein stability prediction Chu et al.
+(2024), protein-protein interaction identification Wang et al.
+(2019); Ofer et al.
+(2021), etc.
 Despite the aforementioned successes, most existing protein language modelling methods suffer from intrinsic limitations.
 Specifically, most of their focuses have primarily been on the amino acid sequence, often neglecting the crucial aspects of protein structure.
-Proteins possess the ability to fold into diverse 3D shapes, interacting with various proteins and small molecules in biologically significant ways jumper2021highly ; mirdita2022colabfold ; tsaban2022harnessing .
-Since protein’s structure determines function greslehner2018molecular , utilizing protein 3D structure information effectively is crucial for protein language modelling, in which many studies have demonstrated the potential of pre-training on experimentally determined protein structures hermosilla2022contrastive ; su2023saprot ; wang2022lm ; zhang2022protein .
+Proteins possess the ability to fold into diverse 3D shapes, interacting with various proteins and small molecules in biologically significant ways Jumper et al.
+(2021); Mirdita et al.
+(2022); Tsaban et al.
+(2022).
+Since protein’s structure determines function Greslehner (2018), utilizing protein 3D structure information effectively is crucial for protein language modelling, in which many studies have demonstrated the potential of pre-training on experimentally determined protein structures Hermosilla and Ropinski (2022); Su et al.
+(2023); Wang et al.
+(2022); Zhang et al.
+(2022).
 Nevertheless, these methods focus only on the structure within proteins and ignore the global similarities between proteins.
 We emphasize that the information on protein structure is not only limited to its structure (i.e., conformation) in 3D space but also includes information ranging from local amino acid molecules to the global structural similarity between proteins, as shown in Figure 1.
-Local information involves the detailed properties and orientations of individual amino acids, which can affect protein stability and biochemical activity renaud2021deeprank .
-These specifics are vital as they demonstrate how modifications or mutations at the amino acid level can alter the overall structure and functionality of the protein jumper2021highly .
-Furthermore, protein structure similarities provide information on evolutionary relationships and functional classes, which are crucial for understanding how structurally similar proteins of different species can perform similar or complementary functions within biological systems hamamsy2023protein .
-For example, as shown in Figure 2, the bacterial ice-binding protein FfIBP and the diatom adhesion protein CaTrailin_4 exhibit no detectable sequence similarity despite their functional similarities zackova2023diatom ; al2023protein .
+Local information involves the detailed properties and orientations of individual amino acids, which can affect protein stability and biochemical activity Renaud et al.
+(2021).
+These specifics are vital as they demonstrate how modifications or mutations at the amino acid level can alter the overall structure and functionality of the protein Jumper et al.
+(2021).
+Furthermore, protein structure similarities provide information on evolutionary relationships and functional classes, which are crucial for understanding how structurally similar proteins of different species can perform similar or complementary functions within biological systems Hamamsy et al.
+(2023).
+For example, as shown in Figure 2, the bacterial ice-binding protein FfIBP and the diatom adhesion protein CaTrailin_4 exhibit no detectable sequence similarity despite their functional similarities Zackova Suchanova et al.
+(2023); Al-Fatlawi et al.
+(2023).
 Their predicted structures exhibit a remarkable similarity (TM-Score = 0.6), with both proteins adopting a beta-helical fold comprised of two units linked by an alpha helix.
 This structural topology is characteristic of ice-binding proteins.
-Such comparisons are key to predicting the functions of newly discovered proteins based on known structures, thereby enhancing our grasp of complex biological processes and interactions lipman1985rapid ; hamamsy2022tm ; hamamsy2023protein .
+Such comparisons are key to predicting the functions of newly discovered proteins based on known structures, thereby enhancing our grasp of complex biological processes and interactions Lipman and Pearson (1985); Hamamsy et al.
+(2022, 2023).
 However, most existing approaches have ineffectively incorporated amino acid molecule information and protein structural similarities into protein representation learning.
@@
 (b) CaTrailin_4
-(c) Structure Alignment
+(c) Alignment
 Figure 2: An example of protein structure similarity.
-Given the predictive structures of a protein pair: (a) the bacterial ice-binding protein FfIBP and (b) the diatom adhesion protein CaTrailin_4 zackova2023diatom ; al2023protein , (c) is FfIBP (blue) and CaTrailin_4 (green) structure alignment.
+Given the predictive structures of a protein pair: (a) the bacterial ice-binding protein FfIBP and (b) the diatom adhesion protein CaTrailin_4 Zackova Suchanova et al.
+(2023); Al-Fatlawi et al.
+(2023), (c) is FfIBP (blue) and CaTrailin_4 (green) structure alignment.
 To eliminate these limitations, we propose a novel protein pre-training framework GLProtein with Global-and-Local Protein structure information for protein representation learning.
@@
 2 Related Work
-In this section, we briefly review related work about protein language modelling as well as protein structure modelling.
-Protein Langauge Modeling.
-As an approach to protein representation learning, protein language modelling is a burgeoning field at the intersection of computational biology and natural language processing (NLP).
-Inspired by the success of language models in NLP, researchers have adapted these techniques to analyze and predict the properties of protein sequences fan2025computational .
-Recent advancements have been dominated by the application of transformer-based models, which utilize self-attention mechanisms to capture relationships between amino acids in a sequence.
-ProtTrans elnaggar2021prottrans and ESM beal2015esm ; verkuil2022language ; hie2022high , trained on large-scale protein databases, have shown remarkable ability in tasks such as protein classification and interaction prediction.
-Moreover, OntoProtein zhangontoprotein and KeAP zhou2023protein incorporated external biological knowledge to enrich protein representations and enhance performance on various downstream tasks.
-However, most of these protein language models do not explicitly consider the spatial structure of proteins and structural similarities between proteins like our proposed approach.
-Figure 3: Overview of our proposed model, which jointly optimizes global protein similarities and masked protein model with local structure information.
+Protein Langauge Modelling.
+As an approach to protein representation learning, protein language modelling is a burgeoning field at the intersection of computational biology and natural language processing (NLP) Hu et al.
+(2023); Li et al.
+(2024b, c, a).
+Inspired by the success of language models in NLP, researchers have adapted these techniques to analyse and predict the properties of protein sequences Fan et al.
+(2025).
+Recent advancements have been dominated by the application of transformer-based models, which utilise self-attention mechanisms to capture relationships between amino acids in a sequence.
+ProtTrans Elnaggar et al.
+(2021) and ESM Beal (2015); Verkuil et al.
+(2022); Hie et al.
+(2022), trained on large-scale protein databases, have shown remarkable ability in tasks such as protein classification and interaction prediction.
+Moreover, OntoProtein Zhang et al.
+and KeAP Zhou et al.
+(2023a) incorporated external biological knowledge to enrich protein representations and enhance performance on various downstream tasks.
+However, most of these protein language models do not explicitly consider the spatial structure of proteins and structural similarities between proteins, like our proposed approach.
+Figure 3: Overview of our proposed model, which jointly optimises global protein similarities and masked protein model with local structure information.
 Protein Structure Modelling.
 The structure of a protein determines its functions.
-Thus, protein structure modelling has been treated as a reliable way to improve protein representation learning huang2024protein ; alquraishi2021machine ; torrisi2020deep ; cheng2008machine .
-Some methods use Graph Neural Networks (GNNs) to handle the complex, non-linear relationships inherent in protein structure jha2022prediction ; reau2023deeprank ; xu2024deeprank .
-Moreover, RGN2 chowdhury2022single utilized a protein language model to learn structural information from unaligned protein sequences.
-GearNet zhang2022protein focused on geometric pertaining and learned protein features by utilizing spatial relationships between amino acids.
-SaProt su2023saprot introduced the concept of a "structure-aware vocabulary" to integrate residue tokens with structure tokens.
-Similar to the knowledge hancing method, PST chen2024endowing enhances protein language models by integrating structural information through graph transformers to incorporate structural data.
-Unlike these models, we propose global structure learning and local structure learning methods, which could not only integrate protein structure information and amino acid information but also learn the structure similarity between different proteins by using TM-Score hamamsy2023protein .
-3 The Proposed Method GLProtein
+Thus, protein structure modelling has been treated as a reliable way to improve protein representation learning Huang et al.
+(2024); AlQuraishi (2021); Torrisi et al.
+(2020); Cheng et al.
+(2008).
+Some methods use Graph Neural Networks (GNNs) to handle the complex, non-linear relationships inherent in protein structure Liao et al.
+(2023); Jha et al.
+(2022); Réau et al.
+(2023); Xu and Bonvin (2024); Zhou et al.
+(2023b, 2024); Li et al.
+(2025).
+Moreover, RGN2 Chowdhury et al.
+(2022) utilized a protein language model to learn structural information from unaligned protein sequences.
+GearNet Zhang et al.
+(2022) focused on geometric pertaining and learned protein features by utilizing spatial relationships between amino acids.
+SaProt Su et al.
+(2023) introduced the concept of a "structure-aware vocabulary" to integrate residue tokens with structure tokens.
+Similar to the knowledge hancing method, PST Chen et al.
+(2024) enhances protein language models by integrating structural information through graph transformers to incorporate structural data.
+Unlike these models, we propose global structure learning and local structure learning methods, which could not only integrate protein structure information and amino acid information but also learn the structure similarity between different proteins by using TM-Score Hamamsy et al.
+(2023).
+3 Methodology
 In this section, we aim to introduce our proposed framework (GLProtein) as a novel solution to learn global and local protein structure information for protein representation learning.
-3.1 An Overview of the Proposed Framework
-Most existing protein language modelling methods often have limitations, primarily concentrating on linear sequence information while overlooking the essential aspects of protein structure.
-To address such limitations, we develop GLProtein, a novel protein pre-training framework that incorporates both global and local protein structure information into protein representation learning.
-More specifically, in order to integrate structure information effectively, a global protein triplet contrastive learning, protein 3D distance encoding, and substructure-based molecular encoding fusion are introduced to empower the protein language model for protein representation learning.
-The framework of GLProtein, shown in Figure 3, consists of three components: protein language modelling, global structure information modelling, and local structure information modelling.
-The first component is protein language modelling, which is the backbone of our proposed framework.
-It employs random masking to the amino acid sequence and treats local structure information as associated structure information, including 3D distance encoding and substructure-based molecular encoding.
-The second component is global structure information modelling, as input data in this component is still protein sequence.
-To obtain a protein triplet, we apply positive and negative protein sampling.
-After that, we use self-supervised contrastive learning to calculate the protein triplet loss and learn the protein similarity with global structure information by taking advantage of the huge amount of unlabeled data.
-The third component is local structure information modelling, which integrates local protein structure information in the protein decoder, including 3D distance encoding and substructure-based molecular encoding.
-Next, we will detail each model component.
+We develop GLProtein that incorporates both global and local protein structure information into protein representation learning.
+The framework of GLProtein, shown in Figure 3, consists of three components: protein language modelling (Section 3.1), global structure information modelling (Section 3.2), and local structure information modelling (Section 3.3).
 3.1 Protein Language Modelling
-Drawing inspiration from the success of masked language modelling in natural language processing, masked protein modelling conceals part of a protein sequence and tasks a model with predicting the hidden residues devlin2018bert ; elnaggar2021prottrans .
 As shown in the center part of Figure  3, protein language modelling forms the backbone of our proposed framework, which aims to learn protein representation.
-We adopt a masking strategy similar to KeAP zhou2023protein , where 20% of the amino acids in a sequence are randomly selected for masking.
-Each masked amino acid has an 80% probability of being masked for prediction, a 10% chance of being replaced with a random amino acid, and a 10% chance of remaining unchanged.
+We adopt a masking strategy that each masked amino acid has an 80% probability of being masked for prediction, a 10% chance of being replaced with a random amino acid, and a 10% chance of remaining unchanged.
 We then integrate protein 3D distance encoding and substructure-based molecular encoding into a protein decoder, in which we will detail in the local structure information modelling component.
@@
 where $\theta_{E}$ and $\theta_{D}$ denote the parameters of the protein sequence encoder and decoder, respectively.
-We initialize with a pre-trained BERT-like encoder: ProtBert elnaggar2021prottrans .
-Next, we will introduce the global structure information modelling part.
+We initialise with a pre-trained BERT-like encoder: ProtBert Elnaggar et al.
+(2021).
 3.2 Global Structure Information Modelling
-In protein science, proteins follow the sequence-structure-function paradigm koehler2023sequence , where the amino acid sequence dictates the protein’s three-dimensional structure, ultimately determining its function.
-Therefore, effectively utilizing information about protein 3D structures is essential for improving protein language modelling.
-However, protein structures encompass more than mere 3D spatial configurations; they also include global structural information that reflects similarities among proteins.
-Understanding these structural similarities is crucial, as they can provide insights into evolutionary relationships and functional roles, allowing for more accurate predictions of protein behaviour and interactions hamamsy2023protein .
-To address this, we introduce the concept of global structure information, which contains the structure similarities between proteins, by leveraging of the huge amount of self-supervised signals in protein sequences, as shown in the top of the Figure 3.
-To be specific, given for each input protein sequence, positive and negative protein sampling is designed to get the triplet $(P,P_{pos},P_{neg})$ for capturing protein structure similarity features.
+Protein structures encompass more than mere 3D spatial configurations; they also include global structural information that reflects similarities among proteins.
+To address this, we introduce the concept of global structure information, which contains the structure similarities between proteins, by leveraging the huge amount of self-supervised signals in protein sequences, as shown at the top of Figure 3.
+To be specific, given each input protein sequence, positive and negative protein sampling is designed to get the triplet $(P,P_{pos},P_{neg})$ for capturing protein structure similarity features.
 Then, the protein triplets are encoded to protein representation for the calculation of the contrastive learning loss.
-This optimizes the protein sequence encoder by bringing the representation of the input protein $P$ and its positive samples $P_{pos}$ closer together while pushing the representation of $P$ and its negative samples $P_{neg}$ further apart in the representation space.
-3.3.1 Positive and Negative Protein Sampling
-TM-score (Template Modeling Score) zhang2004scoring ; xu2010significant is a widely used metric in structural biology for assessing the structural similarity between two protein structures.
-One of its key advantages is the ability to effectively identify similar three-dimensional structures, even among proteins with low sequence similarity.
-Additionally, the TM-score is more sensitive to global fold similarity than to local structural variations, allowing it to capture essential information about the overall architecture of proteins.
-Consequently, we utilize the TM-score to measure structural similarity between proteins, focusing on their overall global structure rather than mere sequence identity.
+This optimises the protein sequence encoder by bringing the representation of the input protein $P$ and its positive samples $P_{pos}$ closer together while pushing the representation of $P$ and its negative samples $P_{neg}$ further apart in the representation space.
+Positive and Negative Protein Sampling.
+TM-score (Template Modeling Score) Zhang and Skolnick (2004); Xu and Zhang (2010) is a widely used metric in structural biology for assessing the structural similarity between two protein structures.
+We utilize the TM-score to measure structural similarity between proteins, focusing on their overall global structure rather than mere sequence identity.
 Mathematically, the TM-score can be expressed as:
-[…]
+$\text{TM-score}=max[\frac{1}{L_{N}}\sum^{L_{r}}_{i=1}\frac{1}{1+(\frac{d_{i}}{d_{0}})^{2}}],$ (2)
 where $L_{N}$ is the length of the native structure, $L_{r}$ is the length of the aligned residues to the template structures, $d_{i}$ is the distance between the $i$-th pair of residues, and $d_{0}$ is a scaling factor.
-We employ a two-pronged approach that utilizes the TM-Vec model hamamsy2023protein to construct a robust set of positive and negative samples for our protein structure similarity analysis.
+We employ a two-pronged approach that utilizes the TM-Vec model Hamamsy et al.
+(2023) to construct a robust set of positive and negative samples for our protein structure similarity analysis.
 For positive sample selection, we utilize the TM-Vec model to identify the top-K protein sequences that exhibit the highest TM-score values in relation to the template proteins.
-This approach ensures that our positive samples share significant structural similarity with the template.
 In contrast, our negative sampling strategy employs a stochastic selection process followed by structural dissimilarity confirmation.
@@
 Subsequently, we employ the TM-Vec model to compute the TM-score between each selected protein and the template protein.
-Proteins with a TM-score < 0.2 are classified as negative samples, as this threshold indicates a high degree of structural dissimilarity xu2010significant .
-This method ensures that our negative samples are structurally distinct from the template, providing a clear contrast to the positive samples.
-3.3.2 Protein Triplet Modelling
+Proteins with a TM-score < 0.2 are classified as negative samples, as this threshold indicates a high degree of structural dissimilarity Xu and Zhang (2010).
+Protein Triplet Modelling.
 After positive and negative protein sampling, we obtain the triplet $(P,P_{pos},P_{neg})$.
@@
 $L_{p}$ denotes the length of amino acid sequence and $D$ stands for the feature dimension.
-To refine the embedding space by ensuring that proteins with similar structures or functions are clustered closer together than those that are dissimilar, we need to specify an objective function to optimize.
 Since the task we focus on in this part is contrastive learning, the protein triplet loss is designed.
@@
 Thus, given protein representation triplet $(P,P_{pos},P_{neg})$, the protein triplet loss $\mathcal{L}_{PTL}$ can be defined as:
-$\begin{split}\mathcal{L}_{PTL}(P,P_{pos},P_{neg})=max(||\mathbf{E}_{p}-\mathbf{E}^{pos}_{p}||_{2}-||\mathbf{E}_{p}-\mathbf{E}^{neg}_{p}||_{2}+\epsilon,0),\end{split}$ (3)
+[…]
 where $\mathbf{E}_{p}$, $\mathbf{E}^{pos}_{p}$, $\mathbf{E}^{neg}_{p}\in\mathbb{R}^{L_{p}\times D}$ are protein representation of the triplet $(P,P_{pos},P_{neg})$.
@@
 3.3 Local Structure Information Modelling
-Complementing global structural information, local protein information offers additional structure-related insights for protein language modelling.
 While the global structure information modelling component is designed to identify structural similarities across different proteins, the local structure information modelling component zooms in on the specific, intricate features of a protein’s internal structure, providing a more nuanced understanding.
 More specifically, in this part, we leverage the local structural details of proteins, including protein 3D distance encoding and substructure-based molecular encoding, to enhance the model’s ability to learn and interpret this local configuration effectively, as shown at the bottom of Figure 3.
-We will first detail the process of protein 3D distance encoding.
-Then, the integration of embedding of amino acid molecules will be illustrated.
-3.4.1 Protein 3D Distance Encoding
-In protein representation learning, the integration of protein 3D coordinates holds immense significance for comprehensively capturing proteins’ structural intricacies and functional behaviours.
-The 3D coordinates provide critical insights into how proteins fold and interact in three-dimensional space, influencing their stability, activity, and specificity liu2022generating ; peng2022pocket2mol ; su2023saprot .
-By incorporating 3D structural information into representation learning models, the model can capture proteins’ geometric and spatial characteristics, which are often overlooked by sequence-based approaches.
-We use AlphaFoldDB [^1] https://alphafold.ebi.ac.uk/ as the 3D protein database and integrate the protein 3D distance encoding ying2021transformers to represent protein 3D structural information to ensure rotational and translational invariance.
-However, proteins are complex macromolecules composed of numerous amino acids, each consisting of multiple atoms.
-Consequently, the atomic coordinate data for each protein is extensive.
-To address this challenge, we propose to take advantage of the alpha-carbon ($\alpha$-C) coordinates rather than the entire protein coordinates in protein representation learning.
-That is because the $\alpha$-C atoms are central to the protein backbone and provide a simplified yet informative representation of the protein’s 3D structure.
+Protein 3D Distance Encoding.
+The 3D coordinates provide critical insights into how proteins fold and interact in three-dimensional space, influencing their stability, activity, and specificity Liu et al.
+(2022); Peng et al.
+(2022); Su et al.
+(2023).
+We use AlphaFoldDB [^1] https://alphafold.ebi.ac.uk/ as the 3D protein database and integrate the protein 3D distance encoding Ying et al.
+(2021) to represent protein 3D structural information to ensure rotational and translational invariance.
+We propose to take advantage of the alpha-carbon ($\alpha$-C) coordinates rather than the entire protein coordinates in protein representation learning.
 By capturing the backbone conformation, $\alpha$-C coordinates effectively convey the protein’s overall shape and folding pattern, which are critical for understanding its function.
 Moreover, leveraging $\alpha$-C coordinates balances capturing essential structural information and maintaining computational efficiency.
-Specifically, the coordinates of each $\alpha$-C are processed to represent the current position of the amino acid in 3D space.
-Then, we encode the Euclidean distance metric to reflect the spatial relation between any pair of amino acids in the 3D space.
-Mathematically, given each amino acid pair $(i,j)$, we first process their Euclidean distance with the Gaussian Basis Kernel function scholkopf1997comparing ,
-$\phi^{k}_{(i,j)}=-\frac{1}{\sqrt{2\pi}|\sigma^{k}|}\exp(-\frac{1}{2}(\frac{\gamma_{(i,j)}||\mathbf{r_{i}}-\mathbf{r_{j}}||+\beta_{(i,j)}-\mu^{k}}{|\sigma^{k}|})^{2}),$ (4)
+[…]
 where $k=1,\dots,K$. $K$ is the number of Gaussian Basis kernels. Then, the 3D distance encoding can be calculated as follows:
@@
 Denote $\Phi^{distance}$ as the matrix form of the 3D distance encoding, whose shape is $n\times n$.
-3.4.2 Substructure-based Molecular Encoding
-Amino acid molecules, as more detailed information about protein localization, play a crucial role in protein representation learning, as they form the essential building blocks of proteins and provide the foundational data for understanding protein structure and function lieu2020amino ; lopez2024biochemistry .
+Substructure-based Molecular Encoding.
+As more detailed information about protein localisation, amino acid molecules play a crucial role in protein representation learning, as they form the essential building blocks of proteins and provide the foundational data for understanding protein structure and function Lieu et al.
+(2020); Lopez and Mohiuddin (2024).
 To learn the fine-grained amino acid structure information, we introduce substructure-based molecular encoding to leverage the inherent relationships between molecule motifs and substructural features in amino acid molecules.
-In practice, we utilize the mol2vec jaeger2018mol2vec to process and derive representations for all amino acid molecules to obtain fine-grained molecular structure information.
-Then we construct substructure-based molecular encoding as fine-grained protein sequence representation, which is the other part of local structure information modelling component.
+In practice, we utilize the mol2vec Jaeger et al.
+(2018) to process and derive representations for all amino acid molecules to obtain fine-grained molecular structure information.
 For protein $P$, we have
-$\mathbf{E}_{a}(P)=\text{Concat}(\mathbf{e}_{x_{1}},\mathbf{e}_{x_{2}},\dots,\mathbf{e}_{x_{i}},\dots,\mathbf{e}_{x_{L}}),\text{where}~{}\mathbf{e}_{x_{i}}\in\mathbb{R}^{1\times d},$ (6)
-$L$ is the length of the protein sequence, $\mathbf{e}_{x_{i}}$ is the $i$-th amino acid molecule embedding, and $d$ stands for the feature dimension of the amino acid molecule.
+$\mathbf{E}_{a}(P)=\text{Concat}(\mathbf{e}_{x_{1}},\mathbf{e}_{x_{2}},\dots,\mathbf{e}_{x_{i}},\dots,\mathbf{e}_{x_{L}}),$
+where $\mathbf{e}_{x_{i}}\in\mathbb{R}^{1\times d}$, $L$ is the length of the protein sequence, $\mathbf{e}_{x_{i}}$ is the $i$-th amino acid molecule embedding, and $d$ stands for the feature dimension of the amino acid molecule.
 3.4 Model Training
@@
 Finally, the pre-training objective of the whole framework will be stated.
-3.5.1 Protein Decoder
-As shown in Figure 3, to enhance the learning of protein representation, we introduce a protein decoder that enables each protein sequence representation to iteratively query and extract relevant information from its associated local structure, including protein 3D distance encoding and substructure-based molecular encoding.
-This process facilitates the integration of critical structural details, aiding in the accurate reconstruction of masked amino acids.
-Specifically, the decoder treats protein representation $\mathbf{E}_{p}$ as a query, while the substructure-based molecular encodings $\mathbf{E}_{a}$ are attended to as keys and values and protein 3D distance encoding $\Phi^{distance}$ is attended to as attention bias.
+Protein Decoder.
+As shown in Figure 3, the decoder treats protein representation $\mathbf{E}_{p}$ as a query, while the substructure-based molecular encodings $\mathbf{E}_{a}$ are attended to as keys and values and protein 3D distance encoding $\Phi^{distance}$ is attended to as attention bias.
 Taking the $i$-th layer as an example, the inputs to the protein decoder include $\mathbf{E}_{p}^{i}$, $\Phi^{distance}$ and $\mathbf{E}_{a}$.
 The substructure-based molecular encoding $\mathbf{E}_{a}$ is firstly queries by $\mathbf{E}_{p}^{i}$ as the key and value:
-$Q_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})W_{Q}^{i},~{}K_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{K}^{i},~{}V_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{V}^{i},$ (7)
-where $W_{Q}^{i}$,$W_{K}^{i}$,$W_{V}^{i}$ are learnable matrices. Norm stands for the layer normalization ba2016layer .
-Then, Attention vaswani2017attention is applied to $\{Q_{p}^{i},K_{a}^{i},V_{a}^{i}\}$, where the representation of protein sequence extract helpful, relevant information from the substructure-based molecular encoding.
+$\begin{split}Q_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})W_{Q}^{i},\\ K_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{K}^{i},\\ V_{a}^{i}=\text{Norm}(\mathbf{E}_{a})W_{V}^{i},\end{split}$
+where $W_{Q}^{i}$$W_{K}^{i}$$W_{V}^{i}$ are learnable matrices. Norm stands for the layer normalization Ba et al. (2016).
+Then, Attention Vaswani et al.
+(2017) is applied to $\{Q_{p}^{i},K_{a}^{i},V_{a}^{i}\}$, where the representation of protein sequence extracts helpful, relevant information from the substructure-based molecular encoding.
 The obtained representation $o_{p}^{i}$ stores the helpful structure information for restoring missing amino aids.
 We then add up $o_{p}^{i}$ and $\mathbf{E}_{p}^{i}$ to integrate information, resulting in the representation $\hat{\mathbf{E}}_{p}^{i}$ as follows:
-$o_{p}^{i}=\text{Attention}(Q_{p}^{i},K_{a}^{i},V_{a}^{i},\Phi^{distance}),~{}~{}\hat{\mathbf{E}}_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})+o_{p}^{i}.$ (8)
+$\begin{split}&o_{p}^{i}=\text{Attention}(Q_{p}^{i},K_{a}^{i},V_{a}^{i},\Phi^{distance}),\\ &\hat{\mathbf{E}}_{p}^{i}=\text{Norm}(\mathbf{E}_{p}^{i})+o_{p}^{i}.\end{split}$
 The resulting representation $\hat{\mathbf{E}}_{p}^{i}$ integrates the helpful, relevant structure information that benefits the restoration of missing amino acids.
 We finally forward $\hat{\mathbf{E}}_{p}^{i}$ through a residual multi-layer perceptron to obtain the output representation of the $i$-th block, which also serves as the input to the $(i+1)$-th block.
-Table 1: Comparisons on amino acid contact prediction.
-seq signifies the distance, measured in terms of amino acid units, between two selected amino acids.
-P@L, P@L/2, P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively.
-The best results are bolded, and the second-best results are underlined.
-$126\leq seq<$ 12 $2412\leq seq<$ 24 $24\leq seq$
+$6\leq seq<12$ $12\leq seq<24$ $24\leq seq$
 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
@@
 GLProtein 0.45(±0.02) 0.55(±0.02) 0.66(±0.01) 0.39(±0.03) 0.48(±0.01) 0.58(±0.02) 0.31(±0.02) 0.40(±0.01) 0.47(±0.02)
-Figure 4: An example of amino acid contacts (top-L predictions for ProteinNet alquraishi2019proteinnet test example TBM-hard#T0912).
+Table 1: Comparisons on amino acid contact prediction.
+seq signifies the distance, measured in terms of amino acid units, between two selected amino acids.
+P@L, P@L/2, P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively.
+The best results are bolded, and the second-best results are underlined.
+Figure 4: An example of amino acid contacts (top-L predictions for ProteinNet AlQuraishi (2019) test example TBM-hard#T0912).
 Raw contact probabilities are shown below the diagonal, top L contacts are shown above the diagonal (blue: true positives, red: false positives, grey: ground-truth contacts).
-By training with both protein 3D distance encoding and substructure-based molecular encoding, the model can not only go beyond simply memorizing sequence patterns but also actively decode the underlying structural and chemical principles that govern protein functionality.
-3.5.2 Pre-training Objective
+Pre-training Objective.
 To estimate the model parameters of GLProtein, we adopt the masked protein modelling object and global protein triplet objective to construct the overall model.
@@
 Figure 5: Results on TAPE Benchmark encompass various evaluations.
-SS is a secondary structure task that is evaluated in CB315.
+$S$ is a secondary structure task that is evaluated in CB315.
 We report medium- and long-range results using P@L/2 metrics in contact prediction task.
 In fluorescence and stability prediction tasks, we use Spearman’s $\rho$ metric for evaluation.
-We also provide a related table in Appendix 7.1.
+We also provide a related table in Appendix A.1.
 4 Experiments
 In this section, we evaluate the generalization ability of the learned protein representation by fine-tuning the pre-trained model across a diverse array of downstream applications, including amino acid contact prediction, protein homology detection, protein stability prediction, protein-protein interaction identification, protein-protein binding affinity prediction and semantic similarity inference.
-4.1 Dataset and Implementation Details
-ProteinKG25  zhangontoprotein offers a comprehensive knowledge graph comprising approximately five million triplets.
-This graph includes nearly 600k proteins, 50k attribute terms and 31 relation terms.
-However, note that we only use the protein sequence data in this dataset.
-Our model does not use the attribute and relation terms extracted from Gene Ontology.
-We used modified pre-training data from KeAP zhou2023protein , which removed overlapping data between ProteinKG25 and downstream datasets.
+Pretraining Datasets.
+Swiss-Prot Boeckmann et al.
+(2003) offers a comprehensive and manually curated protein sequence database that includes nearly 600k protein sequences.
+We use it as pertaining dataset.
 Additionally, we use AlphaFoldDB to obtain the protein 3D coordinate datasets.
+Implementation Details.
 We conducted some experiments and compared GLProtein with baselines regarding pre-training and inference time in contact prediction tasks, as shown in Appendix Table 5.
@@
 During pre-training, GLProtein is trained for 300k steps using a learning rate of 1e-5, weight decay of 0.01 over four GPUs (NVIDIA A6000, 48G Memory each).
+We uniformly fine-tuned all downstream tasks without structural information to ensure fair and unbiased comparisons.
 For the amino acid contact prediction and protein-protein interaction task, we randomly selected five random seeds to fine-tune our model and the baseline model separately and report the results.
-For full implementation details, refer to the provided code.
 4.1 Downstream Tasks
-4.2.1 Amino Acid Contact and Secondary Structure Prediction
-Overview of amino acid contact prediction.
+4.1.1 Amino Acid Contact Prediction
+Overview.
 Amino acid contact prediction is a critical task in computational biology, aiming to identify pairs of amino acids within a protein that are in close spatial proximity.
@@
 The model accomplishes this by generating a probability contact matrix for each input protein.
-We tested the model on the dataset collected and organized by ProteinNet alquraishi2019proteinnet and TAPE rao2019evaluating .
-Overview of secondary structure prediction.
-Secondary structure is a fundamental aspect of computational biology, aiming to determine the local structures of protein segments.
-this task is a sequence-to-sequence task where each input protein is mapped to a type of local structure.
-We report accuracy on a per-amino acid basis on the CB513 dataset cuff1999evaluation .
+We tested the model on the dataset collected and organized by ProteinNet AlQuraishi (2019) and TAPE Rao et al.
+(2019).
 Baselines.
 We evaluate our model compared with ten baselines.
-Specifically, we employed variations of LSTM hochreiter1997long , ResNet he2016deep and Transformer vaswani2017attention proposed by the TAPE benchmark rao2019evaluating .
-ProtBert elnaggar2021prottrans is a BERT-like model pre-trained on UniRef100 suzek2007uniref ; suzek2015uniref .
-ESM-2 rives2021biological ; verkuil2022language ; hie2022high feature a transformer architecture pre-trained on the representative sequences from UniRef50 suzek2007uniref ; suzek2015uniref .
-OntoProtein zhangontoprotein and KeAP zhou2023protein are the most recent knowledge-based pre-training methodologies.
-SaProt su2023saprot is the most recent structure-based protein language model.
-LM-GVP wang2022lm and GearNet zhang2022protein are famous geometric methods for protein representation learning.
+Specifically, we employed variations of LSTM Hochreiter and Schmidhuber (1997), ResNet He et al.
+(2016) and Transformer Vaswani et al.
+(2017) proposed by the TAPE benchmark Rao et al.
+(2019).
+ProtBert Elnaggar et al.
+(2021) is a BERT-like model pre-trained on UniRef100 Suzek et al.
+(2007, 2015).
+ESM-2 Rives et al.
+(2021); Verkuil et al.
+(2022); Hie et al.
+(2022) feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al.
+(2007, 2015).
+OntoProtein Zhang et al.
+and KeAP Zhou et al.
+(2023a) are the most recent knowledge-based pre-training methodologies.
+SaProt Su et al.
+(2023) is the most recent structure-based protein language model.
+LM-GVP Wang et al.
+(2022) and GearNet Zhang et al.
+(2022) are famous geometric methods for protein representation learning.
+We uniformly fine-tuned all downstream tasks without structural information to ensure fair and unbiased comparisons.
+Since all structural tokens are masked, residual information still exists; we substitute Foldseek structure tokens with "#" when fine-tuning SaProt.
 Results.
 Table  1 shows the experimental results of amino acid contact prediction.
-Specifically, we notice that our model GLProtein consistently outperforms other models in short- ($126\leq seq<$ 12), medium- ($2412\leq seq<$ 24) and long-range ($seq>24$) contact predictions.
+Specifically, we notice that our model GLProtein consistently outperforms other models in short- ($6\leq seq<12$), medium- ($12\leq seq<24$) and long-range ($seq>24$) contact predictions.
 Notably, our model demonstrates better performance compared to SaProt, which is also a structure-based language model.
@@
 Figure 4 shows more visually that GLProtein’s prediction on the task of contact prediction is closer to labels, i.e., better performance on long-range contact prediction.
-For the secondary structure (SS-Q3 and SS-Q8), as shown in Figure 5, GLProtein outperforms other baselines in SS-Q8 task and shows competitive performance with ProtBERT, OntoProtein and KeAP in SS-Q3 task.
-Considering the approaches taken by Saprot, LM-GVP, and GearNet, which also incorporate protein structural information, the evident performance superiority of GLProtein over these methods indicates that it offers a more effective option for structure-based protein representation learning.
 We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure.
+More results can be found in the Appendix A.1.
 4.1.2 Protein-Protein Interaction
@@
 The challenge of PPI prediction is approached as a multi-label classification problem.
-We conducted our experiments using three datasets: SHS27K chen2019multifaceted , SHS148K chen2019multifaceted and STRING lv2021learning .
+We conducted our experiments using three datasets: SHS27K Chen et al.
+(2019), SHS148K Chen et al.
+(2019) and STRING Lv et al.
+(2021).
 Both SHS27K and SHS148K are considered subsets of STRING, with proteins excluded if they have fewer than 50 amino acids or exhibit 40% or higher sequence identity.
@@
 Baselines.
-Following OntoProtein zhangontoprotein and KeAPzhou2023protein , we have expanded our baseline models to include four additional methods: DPPI hashemifar2018predicting , DNN-PPI li2018deep , PIPR chen2019multifaceted and GNN-PPI lv2021learning .
-These are incorporated alongside existing baselines such as ProtBert, ESM-2, OntoProtein, KeAP, SaProt, LM-GVP and GearNet, providing a comprehensive set of comparisons in our analysis.
-Table 2: Protein-Protein Interaction Prediction Results.
-Breath-First Search (BFS) and Depth-First Search (DFS) are strategies that split the training and testing PPI datasets.
-The best results are bolded, and the second-best results are underlined.
+Following OntoProtein Zhang et al.
+and KeAPZhou et al.
+(2023a), we have expanded our baseline models to include four additional methods: DPPI Hashemifar et al.
+(2018), DNN-PPI Li et al.
+(2018), PIPR Chen et al.
+(2019) and GNN-PPI Lv et al.
+(2021).
+These are incorporated alongside existing baselines such as ProtBert Elnaggar et al.
+(2021), ESM-2 Beal (2015), OntoProtein Zhang et al.
+, KeAP Zhou et al.
+(2023a), SaProt Su et al.
+(2023), LM-GVP Wang et al.
+(2022), GearNet Zhang et al.
+(2022), DeepInter Lin et al.
+(2023), MAPE-PPI Wu et al.
+(2024), ProLLM Jin et al.
+(2024) and ESM-C Hayes et al.
+(2025), providing a comprehensive set of comparisons in our analysis.
 SHS27k SHS148k STRING
@@
 ProtBert 68.44(±0.78) 72.36(±0.85) 70.06(±0.88) 77.46(±0.62) 66.08(±0.91) 86.45(±0.82)
+DeepInter 77.31(±1.14) 77.18(±0.84) 74.52(±0.99) 76.60(±0.49) 77.82(±0.98) 80.04(±1.18)
 OntoProtein 71.37(±0.84) 76.28(±0.77) 74.60(±0.56) 76.33(±0.69) 75.64(±0.91) 90.23(±0.79)
@@
 […]
+MAPE-PPI 83.63(±0.76) 81.01(±0.58) 84.57(±0.91) 83.62(±0.69) 87.18(±0.82) 87.46(±0.59)
 GearNet 85.46(±0.61) 82.73(±0.69) 80.02(±1.26) 82.28(±0.93) 85.55(±0.74) 88.03(±0.51)
@@
 SaProt 91.18(±0.73) 88.85(±1.04) 90.75(±0.91) 80.67(±0.90) 88.23(±0.81) 88.90(±0.74)
+ProLLM 91.49(±0.91) 88.38(±0.78) 90.90(±1.03) 85.34(±0.61) 87.38(±0.77) 86.99(±0.57)
+ESM-C 92.46(±1.02) 88.14(±0.75) 91.86(±0.51) 86.11(±0.78) 87.41(±0.95) 86.78(±0.66)
 GLProtein 96.32(±0.86) 91.23(±0.92) 93.78(±0.77) 86.14(±0.69) 89.41(±0.66) 91.35(±0.89)
+Table 2: Protein-Protein Interaction Prediction Results.
+Breath-First Search (BFS) and Depth-First Search (DFS) are strategies that split the training and testing PPI datasets.
+The best results are bolded, and the second-best results are underlined.
 Results.
-As shown in Table 2, the results clearly indicate that our method consistently outperforms all other methods, including the structure-based protein language model SaProt, across all datasets and both BFS and DFS evaluation metrics.
+As shown in Table 2, the results clearly indicate that our method consistently outperforms all other methods, including the structure-based protein language model such as SaProt and multimodal protein language model such as ESM-C, across all datasets and both BFS and DFS evaluation metrics.
 The observed decline in performance can be linked to the growing amount of fine-tuning data, transitioning from SHS27k to STRING, which diminished the influence of pre-training.
 We believe that the structural similarities between proteins identified during the pre-training step enable GLProtein to excel in the PPI task, resulting in its outstanding performance.
-4.2.3 Homology Detection, Fluorescence and Stability Prediction
-Overview of homology detection.
-The task of predicting remote homology in proteins can be viewed as a classification problem at the molecular level.
-The objective is to input a protein sequence into the homology detection model, which then identifies the correct types of protein fold.
-In our paper, this presents a significant challenge with 1,195 distinct protein folds to classify.
-We utilize data sources from  hou2018deepsf and present the average accuracy achieved on the fold-level heldout set.
-Overview of fluorescence prediction.
-In the realm of protein science, fluorescence prediction is a vital task that involves estimating the fluorescence properties of proteins.
-This is a regression task where each input protein is mapped to a label measuring the most extreme circumstances in which the protein maintains its fold above a concentration threshold.
-We use the data from  rocklin2017global and use Spearman’s rank correlation coefficient as the metric.
-Overview of stability prediction.
-Stability prediction involves estimating the resilience of a protein’s structure under various environmental conditions, a critical factor in understanding its functional efficacy and therapeutic potential.
-This regression task focuses on predicting the intrinsic stability of proteins, which is essential for assessing their capacity to preserve their structural integrity under severe conditions.
-To assess the effectiveness of our model, we measure its performance using Spearman’s rank correlation coefficient across the entire test set rocklin2017global .
-Baselines. As shown in Figure 5, we included ten protein model as baselines.
-Results.
-As for fluorescence prediction, Figure 5 shows our model has the most competitive performance compared to Transformer and KeAP.
-In the domain of stability prediction, our model again shows the highest performance with a score of 0.81.
-This is significantly higher compared to other models, indicating its potential utility in applications like drug design and protein engineering, where stability is paramount.
-4.2.4 Protein Function Prediction
-Figure 6: Comparisons on semantic similarity inference. The best results are bolded, and the second-best results are underlined.
-Methods MF BP CC Avg
-ESM-2 0.31 0.42 0.28 0.34
-ProtBert 0.41 0.35 0.36 0.37
-OntoProtien 0.41 0.36 0.36 0.38
-SaProt 0.40 0.40 0.39 0.40
-KeAP 0.40 0.40 0.40 0.40
-GLProtein 0.41 0.40 0.39 0.40
-Overview.
-Protein function prediction aims to assign biological or biochemical roles to proteins, and we also regard this task as a sequence classification task.
-Following KeAP zhou2023protein , we divide protein attributes into three groups: molecular function (MF), biological process (BP) and cellular component (CC), and report the Spearman’s rank correlation scores for each group.
-Baselines. We evaluate our model compared with five baselines, including ESM-2, ProtBERT, OntoProtein, SaProt and KeAP.
-Results.
-Table 6 assesses the performance of various computational models in prediction protein functions in three categories: MF, BP and CC.
-Additionally, an average score (Avg) is calculated for each method to provide a holistic view of performance across all categories.
-These models all show a balanced performance in three groups.
-It is worth noting that our model does not use any protein attribute-related knowledge and is comparable to OntoProtein and KeAP, which do.
-It also demonstrates the superiority of our approach.
-We also evaluate our GLProtein with 6 baseline models on protein-protein binding affinity estimation task, which is detailed in Appendix 7.2.
-5 Failure Discussion
+5 Conclusion and Future Work
+In this work, we propose GLProtein, a general protein language model with global and local protein structure information.
+GLProtien outperforms the previous protein representation learning model on most downstream applications, demonstrating the performance superiority of GLProtein.
+In the future, we aim to further enhance GLProtein’s capabilities by exploring novel avenues for incorporating multi-modal data sources, refining the model’s interpretability, and extending its applicability to a wider array of biological contexts.
+6 Limitations
 We have observed that GLProtein underperforms on certain individual tasks.
@@
 We plan to further investigate these issues in our future research.
-6 Conclusion and Future Work
-In this work, we propose GLProtein, a general protein pre-training framework with global and local protein structure information.
-By integrating structure similarities between proteins, protein 3D distance encoding and amino acid molecule embeddings, GLProtien outperforms the previous protein representation learning model on most downstream applications, demonstrating the performance superiority of GLProtein.
-In the future, we aim to further enhance GLProtein’s capabilities by exploring novel avenues for incorporating multi-modal data sources, refining the model’s interpretability, and extending its applicability to a wider array of biological contexts.
+7 Acknowledgements
+The research described in this paper has been partially supported by the National Natural Science Foundation of China (project no.
+62102335 and 62372314), General Research Funds from the Hong Kong Research Grants Council (project no.
+PolyU 15207322, 15200023, 15206024, and 15224524), internal research funds from Hong Kong Polytechnic University (project no.
+P0042693, P0048625, and P0051361).
+This work was supported by computational resources provided by The Centre for Large AI Models (CLAIM) of The Hong Kong Polytechnic University.
+Appendix A Appendix
+Table of Content:
+  • Appendix A.1 Results on TAPE Benchmark
+  • Appendix A.2 Secondary structure prediction
+  • Appendix A.3 Homology Detection, Fluorescence and Stability Prediction
+  • Appendix A.4 Protein Function Prediction
+  • Appendix A.5 Protein-Protein Binding Affinity Estimation
+  • Appendix A.6 Ablation Study
+  • Appendix A.7 Parameter Sensitivity Study
+  • Appendix A.8 Visualization of Protein Representations
+  • Appendix A.9 Time Complexity Analysis
+  • Appendix A.10 Hyper-parameters for Fine-tuning
+A.1 Results on TAPE Benchmark
+In addition to the Figure version, we also provide results on TAPE benchmark in a tabular version.
+As shown in Table 6, our model GLProtein performs competitive performance on many tasks, especially on the contact prediction and stability prediction tasks.
+A.2 Secondary structure prediction
+Overview.
+Secondary structure is a fundamental aspect of computational biology, aiming to determine the local structures of protein segments.
+this task is a sequence-to-sequence task where each input protein is mapped to a type of local structure.
+We report accuracy on a per-amino acid basis on the CB513 dataset Cuff and Barton (1999).
+Baselines.
+We evaluate our model compared with ten baselines.
+Specifically, we employed variations of LSTM Hochreiter and Schmidhuber (1997), ResNet He et al.
+(2016) and Transformer Vaswani et al.
+(2017) proposed by the TAPE benchmark Rao et al.
+(2019).
+ProtBert Elnaggar et al.
+(2021) is a BERT-like model pre-trained on UniRef100 Suzek et al.
+(2007, 2015).
+ESM-2 Rives et al.
+(2021); Verkuil et al.
+(2022); Hie et al.
+(2022) feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al.
+(2007, 2015).
+OntoProtein Zhang et al.
+and KeAP Zhou et al.
+(2023a) are the most recent knowledge-based pre-training methodologies.
+SaProt Su et al.
+(2023) is the most recent structure-based protein language model.
+LM-GVP Wang et al.
+(2022) and GearNet Zhang et al.
+(2022) are famous geometric methods for protein representation learning.
+Results.
+For the secondary structure (SS-Q3 and SS-Q8), as shown in Figure 5, GLProtein outperforms other baselines in SS-Q8 task and shows competitive performance with ProtBERT, OntoProtein and KeAP in SS-Q3 task.
+Considering the approaches taken by Saprot, LM-GVP, and GearNet, which also incorporate protein structural information, the evident performance superiority of GLProtein over these methods indicates that it offers a more effective option for structure-based protein representation learning.
+We attribute the enhancements in performance achieved by GLProtein to its innovative integration of global and local structural information, which allows the pre-trained model to gain a deeper understanding of protein structure.
+A.3 Homology Detection, Fluorescence and Stability Prediction
+Overview of homology detection.
+The task of predicting remote homology in proteins can be viewed as a classification problem at the molecular level.
+The objective is to input a protein sequence into the homology detection model, which then identifies the correct types of protein fold.
+In our paper, this presents a significant challenge with 1,195 distinct protein folds to classify.
+We utilize data sources from  Hou et al.
+(2018) and present the average accuracy achieved on the fold-level heldout set.
+Overview of fluorescence prediction.
+In the realm of protein science, fluorescence prediction is a vital task that involves estimating the fluorescence properties of proteins.
+This is a regression task where each input protein is mapped to a label measuring the most extreme circumstances in which the protein maintains its fold above a concentration threshold.
+We use the data from  Rocklin et al.
+(2017) and use Spearman’s rank correlation coefficient as the metric.
+Overview of stability prediction.
+Stability prediction involves estimating the resilience of a protein’s structure under various environmental conditions, a critical factor in understanding its functional efficacy and therapeutic potential.
+This regression task focuses on predicting the intrinsic stability of proteins, which is essential for assessing their capacity to preserve their structural integrity under severe conditions.
+To assess the effectiveness of our model, we measure its performance using Spearman’s rank correlation coefficient across the entire test set Rocklin et al.
+(2017).
+Baselines. As shown in Figure 5, we included ten protein model as baselines.
+Results.
+As for fluorescence prediction, Figure 5 shows our model has the most competitive performance compared to Transformer and KeAP.
+In the domain of stability prediction, our model again shows the highest performance with a score of 0.81.
+This is significantly higher compared to other models, indicating its potential utility in applications like drug design and protein engineering, where stability is paramount.
+A.4 Protein Function Prediction
+Overview.
+Protein function prediction aims to assign biological or biochemical roles to proteins, and we also regard this task as a sequence classification task.
+Following KeAP Zhou et al.
+(2023a), we divide protein attributes into three groups: molecular function (MF), biological process (BP) and cellular component (CC), and report the Spearman’s rank correlation scores for each group.
+Baselines. We evaluate our model compared with five baselines, including ESM-2, ProtBERT, OntoProtein, SaProt and KeAP.
+Methods MF BP $C$ Avg
+ESM-2 0.31 0.42 0.28 0.34
+ProtBert 0.41 0.35 0.36 0.37
+OntoProtien 0.41 0.36 0.36 0.38
+SaProt 0.40 0.40 0.39 0.40
+KeAP 0.40 0.40 0.40 0.40
+GLProtein 0.41 0.40 0.39 0.40
+Table 3: Comparisons on semantic similarity inference. The best results are bolded, and the second-best results are underlined.
+Results.
+Table 3 assesses the performance of various computational models in prediction protein functions in three categories: MF, BP and CC.
+Additionally, an average score (Avg) is calculated for each method to provide a holistic view of performance across all categories.
+These models all show a balanced performance in three groups.
+It is worth noting that our model does not use any protein attribute-related knowledge and is comparable to OntoProtein and KeAP, which do.
+It also demonstrates the superiority of our approach.
+A.5 Protein-Protein Binding Affinity Estimation
+Overview.
+In this task, we focus on assessing how well protein representations can predict changes in binding affinity caused by protein mutations.
+This regression task involves associating each protein pair with a numerical value.
+Following the methodology described in  Unsal et al.
+(2022), we employ Bayesian ridge regression on the outcomes of element-wise multiplication of representations derived from pre-trained protein models.
+This approach is designed to enhance the accuracy of binding affinity predictions.
+We used the SKEMPI dataset from  Moal and Fernández-Recio (2012) and reported the mean square error of 10-fold cross-validation.
+Baselines.
+We evaluate our model compared with six baselines.
+Specifically, we employed PIPR Chen et al.
+(2019), ProtBert Elnaggar et al.
+(2021), ESM-2 Rives et al.
+(2021); Verkuil et al.
+(2022); Hie et al.
+(2022), SaProt Su et al.
+(2023), OntoProtein Zhang et al.
+and KeAP Zhou et al.
+(2023a).
+PIPR is a siamese-residual-RCNN-based model for multifaceted protein–protein interaction prediction.
+ProtBert is a BERT-like model pre-trained on UniRef100 Suzek et al.
+(2007, 2015).
+ESM-2 feature a transformer architecture pre-trained on the representative sequences from UniRef50 Suzek et al.
+(2007, 2015).
+SaProt is the most recent structure-based protein language model.
+OntoProtein and KeAP are the most recent knowledge-based pre-training methodologies.
+Methods Affinity($\downarrow$)
+PIPR 0.63
+ProtBert 0.58
+ESM-2 0.48
+SaProt 0.58
+OntoProtien 0.59
+KeAP 0.52
+GLProtein 0.52
+Table 4: Comparisons on protein-protein binding affinity prediction, with the best result bolded and the second best underlined.
+The notion $\downarrow$ signifies a preference for lower values, reflecting a superior predictive performance in this context.
+Parameters Resouces Pre-training Inference
+(40 examples)
+ProtBert 400M A single TPU Pod V3-512 400k steps 2.02s
+OntoProtein 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.91s
+KeAP 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.94s
+SaProt 650M 64 NVIDIA 80G A100 GPUs 3 Months 3.02s
+ESM-2 650M - - 2.45s
+GLProtein 400M 4 NVIDIA 48G A6000 GPUs 3 Days (continue pertaining on ProtBert) 1.93s
+Table 5: Comparison of the number of parameters, resources, pre-training time, and inference time for GLProtein and baselines.
+Results.
+Table 4 compares several methods of predicting the binding affinity of protein interactions, where a lower score indicates superior performance.
+GLProtein outperforms PIPR, ProtBert, SaProt and OntoProtein.
+It also shows the competitive performance of KeAP and ESM-2.
+A.6 Ablation Study
+Figure 6: Left: Ablations of three proposed approaches.
+Long-range P@L/2 results are reported for contact prediction.
+Right: Ablations of three proposed approaches.
+F1 scores are reported for protein-protein interaction tasks.
+We investigated the effects of employing diverse protein structure information fusion strategies.
+First of all, the exclusion of the global structure information modelling component (representation as "w/o triplet" in Figure 6) resulted in varying degrees of performance deterioration across contact prediction and protein-protein interaction prediction tasks.
+This observation suggests that our global structure similarities through protein triplet contrastive learning stand out as a more efficacious choice.
+Subsequently, upon removing the proposed substructure-based molecular encoding from the local protein structure information component (denoted as "w/o aa" in Figure 6), we noted a decline in performance by approximately 2.5% and 8% for contact prediction and protein-protein interaction tasks, respectively.
+This underscores the essential role of substructure-based molecular encoding within our proposed methodologies.
+Finally when the protein 3D distance encoding was omitted from the local structure information modelling component (indicated as "w/o coord" in Figure 6), a similar trend of performance degradation was observed, further emphasizing the indispensability of this strategy within our architectural framework.
+A.7 Parameter Sensitivity Study
+In this section, we explore the impact of the parameters in the model on the final performance of our protein model.
+We experimented with the number of protein samples in the protein local structure information modelling component and the coefficient of contrastive learning loss for the protein triplet, respectively.
+As shown in Figure 7, we test the number of protein samples from 1 to 4 on the contact prediction task.
+We observe that as the number of protein samples increases, the performance of our model improves to varying degrees in short-, medium- and long-range contact prediction.
+This also shows that our proposed protein triplet approach indeed enables the protein language model to capture the structural similarity features among proteins.
+Due to computational and memory cost considerations, we ended up constructing 4 protein positive samples and 4 protein negative samples for each protein.
+As shown in Figure 8, we test the value of the coefficient $\alpha$ of contrastive learning loss for the protein triplet.
+We divided the experiment into 6 groups and set the values of $\alpha$ to 0.1, 0.3, 0.5, 1, 3, and 5.
+Then, we evaluated them using the protein-protein interaction prediction task on SHS27k, SHS148k, and STRING datasets, respectively.
+We observe that the model achieves the best performance when the value of $\alpha$ is set to 1.
+Thus, we choose $\alpha=1$ in this paper as our model’s setting.
+Structure Evolutionary Engineering
+$S$-Q3 $S$-Q8 Contact Homology Fluorescence Stability
+SaProt 0.51 0.45 0.37 0.12 0.25 0.46
+LSTM 0.75 0.59 0.26 0.26 0.67 0.69
+Transformer 0.73 0.59 0.25 0.21 0.68 0.73
+ResNet 0.75 0.58 0.25 0.17 0.21 0.73
+ESM-2 0.70 0.54 0.43 0.10 0.30 0.65
+LM-GVP 0.69 0.50 0.43 0.20 0.64 0.69
+GearNet 0.71 0.55 0.44 0.25 0.67 0.78
+ProtBert 0.82 0.67 0.35 0.29 0.61 0.73
+OntoProtein 0.82 0.67 0.40 0.24 0.65 0.74
+KeAP 0.82 0.67 0.45 0.29 0.67 0.75
+GLProtein 0.82 0.68 0.48 0.28 0.67 0.81
+Table 6: Results on TAPE Benchmark.
+$S$ is a secondary structure task that is evaluated in CB315.
+In contact prediction, we test medium- and long-range using P@L/2 metrics.
+In protein engineering tasks, we test fluorescence and stability prediction using Spearman’s $\rho$ metric.
+Figure 7: Parameter sensitivity study on the number of protein samples in the local structure information component.
+Figure 8: Parameter sensitivity study on the value of the coefficient $\alpha$ of contrastive learning loss for the protein triplet in the local structure information component.
+Task Epoch Batch Size Warmup Ratio Learning Rate Freeze Bert Optimizer
+Contact 5 8 0.08 3e-5 False AdamW
+Homology 10 32 0.08 4e-5 False AdamW
+Stability 5 64 0.08 1e-5 False AdamW
+$S$-Q3 5 32 0.08 3e-5 False AdamW
+$S$-Q8 5 32 0.08 3e-5 False AdamW
+Fluorescence 15 64 0.10 1e-3 True Adam
+Table 7: Hyper-parameters for fine-tuning.
+Our Proposed GLProtein
+ESM-2
+KeAP
+ProtBert
+Figure 9: Embedding visualizations of GLProtein, ESM-2, KeAP and ProtBert on SCOPe database.
+A.8 Visualization of Protein Representations
+To facilitate a more intuitive comparison, we utilize t-SNE to visualize the protein representations produced by GLProtein, ESM2, KeAP, and ProtBert.
+The visualization results, based on the non-redundant subset ($PIDE\leq 40\%$) of the SCOPe database Chandonia et al.
+(2019), are illustrated in Figure 9.
+As depicted in this figure, the representations for alpha and beta proteins generated by GLProtein are distinctly separated, whereas those produced by ESM-2, KeAP, and ProtBert are more closely intertwined.
+A.9 Time Complexity Analysis
+We provide a more specific complexity analysis as follows: protein encoder operates at approximately $O(L^{2}d)$, where $L$ is the length of protein sequence and $d$ is the embedding dimension.
+Triplet protein sampling operates at approximately $O(L^{3})$, reducing the complexity to $O(L^{2})$ by TM-Vec.
+Triplet loss operates at approximately $O(3Ld)\rightarrow O(Ld)$.
+Protein 3D distance encoding operated at approximately $O(KL^{2}d)$, where $K$ is the number of Gaussian Basis kernels.
+Substructure-based molecular encoding operates at approximately $O(Ld)$.
+Protein decoder operates at approximately $O(L^{2}d)$.
+Total computation cost operated at $O_{total}=O_{encoder}+O_{global}+O_{local}+O_{decoder}=O(L^{2}d)+O(L^{2})+O(Ld)+O(KL^{2}d)+O(L^{2}d)=O((K+1)L^{2}d)$.
+A.10 Hyper-parameters for Fine-tuning
+The hyper-parameters for fine-tuning are provided in the Table 7.
+Specifically, we follow the hyper-parameter settings in GNN-PPI Lv et al.
+(2021) for PPI prediction.
+For protein binding affinity prediction and semantic similarity inference, we follow the fine-tuning configuration in PROBE Unsal et al.
+(2022).